### v2

# From Bugs to Benchmarks: A Comprehensive Survey of Software Defect Datasets

Abstract.
Software defect datasets, which are collections of software bugs, are essential resources to facilitate empirical research and enable standardized benchmarking for a wide range of software engineering techniques, including emerging areas like agentic AI-based software development. Over the years, numerous software defect datasets have been developed, providing rich resources for the community, yet making it increasingly difficult to navigate the landscape. This article provides a comprehensive survey of 151 software defect datasets, covering their scope, construction, availability, usability, and practical uses. We also suggest potential opportunities for future research based on our findings, such as addressing underrepresented kinds of defects.
software bugs, software quality and reliability, benchmarking, empirical software engineering
This work was supported by the National Science Foundation (awards CNS-2016735 and CCF-2119348), by the European Research Council (ERC, grant agreements 851895 and 101155832), and by the German Research Foundation through the ConcSys, DeMoCo, and QPTest projects.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXcopyright: nonejournal: CSURccs: Software and its engineering Software defect analysisccs: Software and its engineering Software reliabilityccs: Software and its engineering Empirical software validationccs: General and reference Surveys and overviewsccs: General and reference Empirical studiesccs: General and reference Evaluation
1. Introduction
Given the escalating scale and complexity of real-world software systems, software defects are inevitable and have become increasingly challenging to manage (Fenton and Ohlsson, 2000; Chou et al., 2001). Software defects, commonly referred to as bugs, are unintentional errors in software that can cause software systems to behave unexpectedly or fail to meet their intended requirements. To address the challenges posed by software defects, in the last few decades, a variety of techniques have been proposed to detect, localize, and fix them. These techniques include, but are not limited to, static and dynamic analysis (Ayewah et al., 2008; Ball, 1999), test generation and prioritization (Nebut et al., 2006; Elbaum et al., 2002), fault detection and localization (Wong et al., 2016), and automated program repair (Le Goues et al., 2021). To evaluate the effectiveness and efficiency of these techniques, the research community heavily relies on software defect datasets, which serve as foundational resources for benchmarking and validation.
Software defect datasets are curated collections of real-world software bugs. These datasets serve as standardized benchmarks, fostering consistent and reproducible evaluations of novel techniques across diverse studies. Additionally, the realism, diversity, scale, and accessibility of these datasets support empirical analyses of defect characteristics and enable the training and fine-tuning of machine learning models for software quality assurance. For instance, the widely adopted Defects4J dataset (DBLP:conf/issta/JustJE14) has been used for a wide range of technical evaluations (e.g., (Martinez et al., 2017; Zhong et al., 2025; Ouyang et al., 2024)), empirical studies (e.g., (Sobreira et al., 2018; Durieux et al., 2019)), and to train machine learning models (e.g., (Zhang et al., 2024b)). As another example, the recently proposed SWE-Bench (DBLP:conf/iclr/JimenezYWYPPN24) dataset is widely used to evaluate the effectiveness of large language models (LLMs) in automating software engineering (Zhang et al., 2024c; Yang et al., 2024) and serves as the basis for other benchmarks (Aleithan et al., 2024; Mündler et al., 2024), demonstrating the practical value of software defect datasets.
With the continuous growth in the complexity and scale of modern software systems, there is an increasing demand for comprehensive and high-quality defect datasets. This demand is further amplified by the rising popularity of artificial intelligence (AI)-based software engineering approaches (Hou et al., 2024; Yang et al., 2022). In response to this escalating demand, industry challenges, competitions, and workshops have been organized to encourage the development of software defect datasets of high quality. As a result, these datasets have experienced a significant increase in both quantity and diversity in recent years, becoming essential resources for researchers and practitioners in multiple fields, including software engineering, programming languages, computer security, computer systems, and artificial intelligence.
With the increasing number of software defect datasets, researchers and practitioners face challenges in navigating the landscape of software defect datasets for two main reasons. First, different datasets may have different focuses, such as types of defects or programming languages, requiring additional effort to identify the most suitable dataset for specific research or practical needs. Second, based on how the datasets are constructed, they exhibit diverse characteristics, such as size and usability, affecting the impact of these datasets. Moreover, creators and maintainers of software defect datasets also encounter challenges in developing datasets that address the evolving needs of software engineering research and practice, as well as in ensuring the long-term sustainability and usability of the created datasets.
Figure 1. Topics covered in this survey.
To address these challenges, this paper provides a comprehensive overview of existing software defect datasets. Specifically, we follow a systematic methodology (detailed in Section 2) for collecting and filtering relevant literature, resulting in the identification of 151 software defect datasets. Our survey focuses on the following four research questions (RQs), as illustrated in Figure 1:
  • RQ1: What is the scope of existing software defect datasets? (Section 3) For each dataset surveyed, we examine the application domain of the software systems from which the defects are collected (e.g., machine learning systems or mobile applications), the types of defects (e.g., functional, security, or performance), the programming languages (e.g., Java, C/C++, or Python), and other characteristics that define the focus of the dataset. We also explore potential interrelationships among these dimensions to provide a holistic understanding of the dataset landscape.
  • RQ2: How are existing software defect datasets constructed? (Section 4) For each dataset, we investigate how the dataset is constructed. This includes the sources of defects (e.g., issue reports, version control systems, or continuous integration & delivery) and the methodologies used to obtain the defects (e.g., manual inspection, keyword matching, or machine learning techniques).
  • RQ3: To what extent are existing software defect datasets available and usable in practice? (Section 5) For each dataset, we evaluate its availability and usability. Availability refers to whether the dataset is publicly available, and if so, where it is hosted. Usability refers to how the defects are presented in the dataset, such as providing only textual descriptions, code snippets, or executable artifacts to reproduce the defects.
  • RQ4: How are software defect datasets actually used by research community? (Section 6) While prior literature reviews (Hirsch and Hofer, 2022; Holek et al., 2023) have provided high-level summaries of the intended uses of software defect datasets from the perspective of their creators, this research question seeks to offer a more comprehensive and detailed understanding of how these datasets are actually used. By analyzing citations of the most influential datasets, we aim to uncover usage patterns and trends over time, thereby revealing their impact on the research community.
Finally, based on our analysis of existing datasets, Section 7 identifies potential opportunities for future research, including automating the mining of high-quality datasets, e.g., by leveraging CI pipelines and LLMs. Additionally, we identify the need for specialized datasets targeting specific domains, defect types, and new languages, as well as techniques ensuring dataset reproducibility and continuous updates.
Our work contributes the first comprehensive survey on software defect datasets, providing a detailed discussion of their scope, construction, availability, and real-world usage. Orthogonal to existing surveys on software defects, which focus on categorizing (Nagwani and Verma, 2014; Zhang et al., 2015), detecting/localizing (Wong et al., 2016; Lin et al., 2020; Paltenghi and Pradel, 2024; Chen et al., 2021; Andreasen et al., 2017), and repairing software defects (Monperrus, 2018; Huang et al., 2023; Zhang et al., 2024a), this survey focuses on datasets that collect software defects prevalent in software codebases. Previous literature reviews on 73 benchmarks (Hirsch and Hofer, 2022) and 14 bug datasets (Holek et al., 2023) provide high-level overviews of existing benchmarking datasets. However, as summarized in Table 1, these studies do not cover several important aspects of software defect datasets. In contrast, this article provides (1) significantly broader coverage of 151 software defect datasets with a more recent cutoff date, (2) a systematic and in-depth analysis of their scope, construction, availability, and usability, (3) insights into the relationships between different dataset characteristics, and (4) an examination of longitudinal trends in the actual usage of the most influential datasets.
We envision this survey to be useful for both dataset users and creators. For dataset users, this survey provides a comprehensive overview of existing software defect datasets, enabling them to identify the most suitable datasets for their research or practical needs. For dataset creators, our work highlights the current focuses and methodologies used in existing datasets, which can guide the development of new datasets that fill the gaps in the current literature. In addition to this article itself, we provide an interactive website [^1] https://defect-datasets.github.io/ that allows users to find datasets based on specific criteria, such as programming languages or application domains of interest.
Table 1. Comparison between surveys of software defect datasets.
Hirsch and Hofer (2022) Holek et al. (2023) This Survey
Number of Datasets 73 14 151
Cutoff Date For Dataset Selection November 2021 April 2022 January 2025
Scope of Datasets
- Application Domains
- Types of Defects
- Programming Languages
Dataset Construction
- Source of Defects
- Mining Methodology
Availability and Usability
- Availability and Hosting Platforms
- Presentation
Uses of Datasets
2. Survey Methodology
Table 2. Searching strategies to collect papers
Strategy Search Content
Key Phrase bug benchmark, bug dataset, defect benchmark, defect dataset, patch benchmark, patch dataset, fail pass, fault benchmark, fault dataset, mining software, software benchmark, software bug, software dataset, software defect, software fault, software patch, software vulnerabilit(y/ies), vulnerabilit(y/ies) benchmark, vulnerabilit(y/ies) dataset, empirical bug, empirical defect, empirical error, empirical failure, empirical fault, empirical issue, empirical patch, study bug, study defect, study error, study failure, study fault, study issue, study patch
Keyword benchmark, bug, dataset, defect, fault, issue, patch, vulnerabilit(y/ies)
Venue ASE, ESE, FSE, ICSE, ISSRE, ISSTA, MSR, OOPSLA, SANER, TOSEM, TSE, ESEC, AAAI, ACL, EMNLP, ICLR, ICML, NeurIPS, USENIX Security, IEEE S&P, CCS, NDSS
Figure 2. Collection and selection process for software defect dataset papers.
This section describes the methodology used in this survey to identify, select, and analyze existing software defect datasets, guided by research questions described in the introduction. Our survey covers 151 software defect datasets. To identify these datasets, we follow a systematic process that collects papers from multiple bibliographic databases and selects papers through a rigorous manual review. Figure 2 illustrates the complete paper collection and selection process. We conduct a comprehensive literature search across four major bibliographic databases: ACM Digital Library, DBLP, IEEE Xplore, and ScienceDirect. For each source, we execute two complementary search strategies, summarized in Table 2. We include papers that have been peer-reviewed and published before January 1, 2025, excluding preprints.
(1) Key Phrase-Based Search
We use specific key phrases to identify papers explicitly related to software defect datasets or empirical studies on software defects. We search for papers containing all words from a key phrase in their titles, abstracts, and keywords if available. This strategy yields 823 papers from the ACM Digital Library, 10,866 papers from DBLP, 9,107 papers from IEEE Xplore, and 276 papers from ScienceDirect.
(2) Keyword ×Venue Search
To complement the key phrase-based search, we use individual keywords combined with venue restrictions. As individual keywords yield many papers, we restrict this search to high-impact venues in software engineering, machine learning, and software security. We search for papers containing at least one keyword in their title, abstract, and keywords if available. This strategy yields 431 papers from the ACM Digital Library, 6,402 papers from DBLP, and 1,299 papers from IEEE Xplore. Due to API query constraints, this strategy was not applicable to ScienceDirect.
After merging results from both search strategies across all bibliographic databases and performing cross-source deduplication, we obtain 23,057 unique papers potentially related to software defect datasets. Two authors independently review the titles and abstracts of these papers to identify those potentially introducing a software defect dataset. The authors then compare their selections and reach consensus through discussion, excluding 22,728 irrelevant papers and identifying 329 papers for full-text assessment. For the full-text assessment phase, the same two authors independently review each of the 329 papers to determine their eligibility for inclusion. A paper is included if its main contribution is one or more software defect datasets, or if it includes software defects as byproducts. Papers are excluded if they only use existing datasets without introducing a new dataset, present datasets for purposes unrelated to software defects (e.g., code generation or performance benchmarking), or focus solely on techniques for mining, injecting, or labeling defects. The inter-rater agreement between the two authors for the full-text assessment is almost perfect, with a Cohen’s kappa (Cohen, 1960) of 0.871 (95% CI: 0.818–0.924). The authors agreed on 308 out of 329 papers (93.6%), and disagreements on the remaining 21 papers were resolved through discussion and, when necessary, consultation with a third author. This rigorous assessment ultimately leads to 150 papers, from which we identified 151 software defect datasets (one paper (DBLP:journals/tse/GouesHSBDFW15) presents two datasets).
Finally, the same two authors conduct an in-depth manual review of all included datasets to extract and annotate their attributes (e.g., application domain, defect type, and programming language for RQ1; sources of defects and construction methodologies for RQ2; and availability and usability for RQ3). Each dataset is independently cross-checked by both authors to ensure consistency, with ambiguities resolved through discussion. The annotation process is performed manually over several months.
(a) Number of papers by year.
(b) Number of papers by venue.
Figure 3. General statistics of software defect dataset papers.
Figure 3(a) plots the number of papers related to software defect datasets published by year, which shows a significant increase in the number of new software defect datasets published in the past eight years. The venues that publish papers related to software defect datasets are shown in Figure 3(b), where different tracks of a conference are grouped together. The top venues that publish papers related to software defect datasets include MSR, ICSE, and ASE. All venues that have only one paper included in this study are grouped into the others category, such as ICLR and SoCC.
3. Scope of Software Defect Datasets (RQ1)
Software defect datasets often have specific purposes. In this section, we discuss and compare the scope of the various datasets. In particular, we consider three dimensions: (1) the application domain of the software that contains the defects (Section 3.1), (2) the type of defects that are included in the dataset (Section 3.2), and (3) the programming languages used in the defective software (Section 3.3).
3.1. Application Domains of Defective Software
Table 3. Application domains of defective software. One dataset may appear in multiple categories.
Category Datasets Count
Machine Learning Applications (DBLP:journals/ese/MorovatiNKJ23; DBLP:conf/icse/LiangLSSFD22; DBLP:conf/msr/KimKL21; DBLP:conf/issta/ZhangCCXZ18; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/kbse/WangWCCY22; DBLP:conf/icse/HumbatovaJBR0T20) 7
Frameworks (DBLP:conf/icse/GuanXLLB23; DBLP:conf/issre/LiuZDHDM022; DBLP:conf/icsm/HoMI0SKNR23; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:journals/ese/TambonNAKA24) 5
Compilers (DBLP:conf/issre/Du0MZ21; DBLP:conf/sigsoft/ShenM0TCC21) 2
Compilers & Runtime Engines (DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/qrs/YuW24; DBLP:conf/icse/PaulTB21a; DBLP:conf/issre/Du0MZ21; DBLP:conf/msr/ZamanAH12; DBLP:conf/uss/XuLDDLWPM23; DBLP:journals/infsof/WangBWYGS23; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/kbse/RomanoLK021) 11
Mobile Applications (DBLP:conf/pst/TebibAAG24; DBLP:conf/kbse/DasAM24; DBLP:conf/secrypt/SenanayakeKAP023; DBLP:conf/issta/XiongX0SWWP0023; DBLP:conf/msr/WendlandSMMHMRF21; DBLP:conf/msr/RiganelliMMM19; DBLP:conf/promise/MitraR17; DBLP:journals/ese/LiuWWXCWYZ19; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/icse/TanDGR18) 10
Blockchain Smart Contracts (DBLP:journals/jss/CaiLZZS24; DBLP:journals/tse/ChenXLGLC22; DBLP:conf/icsm/ZhangXL20; DBLP:journals/jss/WangCHZBZ23) 4
Clients (DBLP:conf/msr/KimKL22) 1
Cloud Computing (DBLP:conf/issta/XuG024; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/sigsoft/GaoDQGW0HZW18; DBLP:journals/pacmpl/DrososSAM024) 4
Autonomous Systems (DBLP:journals/ese/TimperleyHSDW24; DBLP:conf/icse/GarciaF0AXC20; DBLP:conf/sigsoft/WangLX0S21) 3
Quantum Computing Applications (DBLP:journals/jss/ZhaoMLZ23) 1
Platforms (DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/qsw/ZhaoWLLZ23) 2
Scientific Computing (DBLP:conf/hotsos/MurphyBSR20; DBLP:conf/msr/AzadIHR23; DBLP:conf/kbse/FrancoGR17) 3
Databases (DBLP:journals/tosem/LiuMC24; DBLP:conf/icse/CuiD0WSZWYXH0024) 2
Computational Notebooks (DBLP:journals/tosem/SantanaNAA24) 1
Internet of Things (DBLP:conf/icse/Makhshari021) 1
For clarity, only datasets that explicitly specify the corresponding attribute (i.e., application domain) are listed. The complete list of 151 datasets is available at https://defect-datasets.github.io/ and in the supplementary materials (Zhu et al., 2025). The same convention applies to all tables in Sections 3 and 4.
The application domain of defective software refers to the specific field or area in which the software operates, such as machine learning or cloud computing. Of the 151 software defect datasets considered here, the majority (over $60\%$) focus on general software development and are not tied to any specific application domain. Among the remaining datasets that explicitly target specific domains, we identify eleven distinct application domains (Table 3), including machine learning, compilers and runtime engines, mobile applications, and blockchain. [^2] In this and other classifications of datasets, a single dataset may appear in multiple categories. Some domains attract more attention from dataset creators than others. For example, there are 14 datasets related to machine learning defects, which we further categorize into machine learning applications, machine learning frameworks, and machine learning compilers. The rest of this section discusses the application domains focused on by the software defect datasets, highlighting the uniqueness of defects in each domain.
3.1.1. Defects Related to Machine Learning
The reliability and robustness of machine learning applications and their underlying software platforms are crucial for their successful deployment. Machine learning-related software defects can cause severe consequences, such as incorrect predictions, biased decisions, and performance degradation (DBLP:conf/issta/ZhangCCXZ18). We identify three subdomains of machine learning-related software, each with several defect datasets: machine learning applications, machine learning frameworks, and machine learning compilers.
Machine Learning Applications
Defects in machine learning applications often arise during model development, training, and deployment, leading to critical issues like incorrect predictions or silent errors. Datasets such as defect4ML (DBLP:journals/ese/MorovatiNKJ23), gDefects4DL (DBLP:conf/icse/LiangLSSFD22), and Denchmark (DBLP:conf/msr/KimKL21) focus on common machine learning defects, including API misuses, performance problems, and numerical errors. There are also machine learning defect datasets provided as by-products of empirical studies, such as studies on bugs from programs relying on TensorFlow (DBLP:conf/issta/ZhangCCXZ18), on numerical defects in deep learning programs (DBLP:conf/kbse/WangWCCY22), and on deployment-related bugs in machine learning-based mobile applications (DBLP:conf/icse/ChenYLCLWL21).
Machine Learning Frameworks
Machine learning frameworks, such as PyTorch and TensorFlow, are crucial for building and optimizing models, but themselves may suffer from defects. Software defect datasets for machine learning frameworks typically result from empirical research. For example, MOB-Dataset (DBLP:conf/icse/GuanXLLB23) collects real-world machine learning optimization defects from TensorFlow and PyTorch, and comes with an empirical study of model optimization. Likewise, empirical studies of machine learning frameworks have produced defect datasets of common bugs in PyTorch (DBLP:conf/icsm/HoMI0SKNR23) and quantum machine learning frameworks (DBLP:conf/qsw/ZhaoWLLZ23). Other work presents a dataset and a defect taxonomy of aging-related bugs in machine learning libraries (DBLP:conf/issre/LiuZDHDM022). These studies provide critical insights into the defects that can impact the functionality, efficiency, and stability of machine learning frameworks.
Machine Learning Compilers
In addition to applications and frameworks, machine learning compilers play a vital role in optimizing models, but they also introduce unique bugs. Based on datasets of defects in machine learning compilers, DBLP:conf/issre/Du0MZ21 and DBLP:conf/sigsoft/ShenM0TCC21 both examine frequent issues in these compilers, such as incorrect optimizations, numerical inaccuracies, and incompatibilities with hardware accelerators.
3.1.2. Defects in Compilers and Runtime Engines
Compilers and runtime engines are critical components of software systems, responsible for translating and executing code. Defects in compilers and runtime engines can significantly impact software performance, security, and execution correctness. We here mean by “runtime engine” both stand-alone implementations of a programming language, e.g., the Java Virtual Machine, and engines integrated into a larger software, e.g., a JavaScript engine embedded into a web browser. Datasets of defects in compilers and runtime engines often come as by-products of empirical research. For example, DBLP:conf/qrs/YuW24 conduct an empirical study on bugs in implementations of the Rust language , presenting a dataset of 194 bugs from Rust compilers and Rust infrastructure. DBLP:conf/uss/XuLDDLWPM23 investigate security bugs introduced by compilers for C/C++, highlighting the potential risks when silent bugs go undetected. In addition, as also mentioned above, DBLP:conf/issre/Du0MZ21 and DBLP:conf/sigsoft/ShenM0TCC21 study defects in deep-learning compilers, identifying common root causes and symptoms. For runtime engines, Bugs in Pods (DBLP:conf/issta/YuXZ0LS24) provides 429 bugs in the container runtime systems containerd , cri-o , gvisor , and runc . Additionally, DBLP:conf/icse/PaulTB21a provide a detailed dataset of security-related changes in the Chromium OS project, and DBLP:conf/msr/ZamanAH12 present a dataset of performance-related bugs in web browsers, focusing on how such bugs degrade system performance. Besides browsers, datasets constructed from studies of other runtime engines include defects in WebAssembly engines (DBLP:conf/wcre/WangZRLJ23; DBLP:journals/tosem/ZhangCWCLMMHL24), and bugs that impact the correctness and performance of JavaScript engines (DBLP:journals/infsof/WangBWYGS23). Altogether, these datasets are critical for understanding and addressing defects in compilers and runtime environments.
3.1.3. Defects in Mobile Applications
Mobile applications are prone to various types of defects, including functional bugs, security vulnerabilities, and resource issues. Several datasets, mostly based on Android applications, have been developed to represent these defects. For example, LVDAndro (DBLP:conf/secrypt/SenanayakeKAP023) and Ghera (DBLP:conf/promise/MitraR17) focus on security vulnerabilities, providing labelled vulnerability datasets and benchmarks for AI-based detection models and vulnerability analysis. Functional bugs are collected by Andror2 (DBLP:conf/msr/WendlandSMMHMRF21) and by DBLP:conf/issta/XiongX0SWWP0023, both of which analyze the behavior of defective mobile applications by manually reproducing the bugs. More specialized defects, such as data loss issues and resource leaks, are captured by DataLossRepository (DBLP:conf/msr/RiganelliMMM19) and DroidLeaks (DBLP:journals/ese/LiuWWXCWYZ19), respectively. Finally, DroixBench (DBLP:conf/icse/TanDGR18) provides defective APK files particularly for evaluating automated repair techniques on Android applications. These datasets play a crucial role in enhancing the reliability and security of Android applications by supporting bug detection, repair, and vulnerability research.
3.1.4. Defects Related to Blockchain
Blockchain, as a decentralized and distributed ledger technology, has been widely used in various applications, such as cryptocurrencies (Zheng et al., 2017). In general, blockchain applications are implemented as smart contracts, which are self-executing contracts with the terms of the agreement between multiple parties directly written into code. Defects in blockchain applications can have severe implications, such as financial loss. Several datasets focus on identifying and analyzing defects in smart contracts. From a general perspective, Jiuzhou (DBLP:conf/icsm/ZhangXL20) is a dataset of Ethereum smart contract bugs that are classified into 49 categories. Specifically for security vulnerabilities, SBP-Dataset (DBLP:journals/jss/CaiLZZS24) contains smart contract defects generated by removing guards and turning safe functions into unsafe ones. In addition, there are two datasets constructed from empirical studies. One defines 20 categories of smart contract bugs (DBLP:journals/tse/ChenXLGLC22), and the other identifies common patterns of bug fixes (DBLP:journals/jss/WangCHZBZ23). Unlike the above datasets, DBLP:conf/msr/KimKL22 focus on energy bugs in Ethereum client software instead of smart contracts. The dataset contains 507 bugs that may cause energy inefficiency mined from the buggy commits of client software operated in the Ethereum network.
3.1.5. Defects in Cloud Computing
Cloud computing systems provide on-demand access to computing resources over the internet and are prone to various software defects that can affect performance, reliability, and security. Datasets of software defects in cloud computing often arise from empirical studies. For instance, both DBLP:conf/cloud/GunawiHLPDAELLM14 and DBLP:conf/sigsoft/GaoDQGW0HZW18 investigate real-world issues in distributed cloud systems, with the latter focusing on crash recovery bugs. Additionally, DBLP:conf/issta/XuG024 present a dataset of bugs in cloud orchestration systems, while DBLP:journals/pacmpl/DrososSAM024 provide a dataset of misconfigurations in cloud infrastructure-as-code scripts.
3.1.6. Defects in Autonomous Systems
Autonomous systems, such as autonomous vehicles, aircraft, or robots, rely on robust software to operate safely and efficiently. Defects in autonomous systems can lead to severe consequences, such as accidents, injuries, and fatalities. DBLP:conf/sigsoft/WangLX0S21 and DBLP:conf/icse/GarciaF0AXC20 both present empirical studies of software defects in autonomous systems. The former study has collected a dataset of 569 bugs from PX4 and ArduPilot , which are well-known open-source autopilot software systems for unmanned aerial vehicles. The latter study presents a dataset of 499 autonomous vehicle bugs from Apollo and Autoware . In addition, ROBUST (DBLP:journals/ese/TimperleyHSDW24) offers a dataset of 221 bugs from seven popular robot operating systems. Together, these datasets highlight the unique challenges caused by defects in autonomous systems.
3.1.7. Defects Related to Quantum Computing
Defects in quantum computing systems, including both quantum applications and quantum platforms, present unique challenges due to the complexity and novelty of the technology. Bugs4Q (DBLP:journals/jss/ZhaoMLZ23) provides a benchmark of bugs in quantum programs, particularly focusing on the Qiskit framework , enabling controlled testing and debugging studies. Additionally, there are defect datasets constructed within empirical studies, such as one that gathers and categorizes bug patterns across various quantum platforms (DBLP:journals/pacmpl/PaltenghiP22), and another that focuses on defects within the intersection of quantum computing and machine learning (DBLP:conf/qsw/ZhaoWLLZ23). These datasets are useful for advancing the reliability of quantum software systems.
3.1.8. Defects Related to Scientific Computing
Scientific computing systems are often complex and require high performance and accuracy. Defects in scientific computing can lead to incorrect results, degraded performance, or even compromised system security. DBLP:conf/hotsos/MurphyBSR20 collect security defects from open-source scientific software projects written in Julia, a programming language designed for numerical analysis and computational science. Likewise, DBLP:conf/msr/AzadIHR23 build a dataset by investigating 1,729 performance bugs from 23 real-world high-performance computing (HPC) applications such as molecular simulation and computational fluid dynamics. In addition to scientific computing applications, DBLP:conf/kbse/FrancoGR17 study numerical bugs across scientific computing libraries and present a dataset of 269 numerical bugs from six libraries, including NumPy , SciPy , and LAPACK . Collectively, these datasets provide valuable insights into the functional, performance, and security challenges faced in scientific computing.
3.1.9. Other Application Domains
In addition to the above, software defect datasets have been created for various other application domains. For example, for database-related software, datasets are available for both transaction bugs in database systems (DBLP:conf/icse/CuiD0WSZWYXH0024) and access bugs in database applications (DBLP:journals/tosem/LiuMC24). Other application domains include computational notebooks (DBLP:journals/tosem/SantanaNAA24) and internet of things (IoT) (DBLP:conf/icse/Makhshari021). While these domains have only a small number of datasets, these datasets provide insights into the challenges caused by domain-specific bugs.
3.2. Types of Defects
Table 4. Types of defects. One dataset may contain multiple types.
Type Datasets Count
Functional (DBLP:journals/ese/TimperleyHSDW24; DBLP:conf/edcc/AndradeLV24; DBLP:conf/msr/LiuHLZCSHM24; DBLP:journals/tosem/LiuMC24; DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:conf/icse/CuiD0WSZWYXH0024; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/issta/XuG024; DBLP:conf/ease/WaseemDA0M24; DBLP:conf/qrs/YuW24; DBLP:conf/acl/TianYQCLPWHL0024; DBLP:conf/iclr/JimenezYWYPPN24; DBLP:conf/msr/PramodSTSW24; DBLP:conf/msr/SilvaSM24; DBLP:conf/kbse/SongWCCLLWP23; DBLP:conf/icst/LeeKYY24; DBLP:conf/kbse/AnKCYY23; DBLP:conf/sigsoft/WuLZ024; DBLP:conf/kbse/AvulaVM23; DBLP:conf/msr/MahbubSR23; DBLP:journals/jss/ZhaoMLZ23; DBLP:conf/icst/KimH23; DBLP:journals/ese/MorovatiNKJ23; DBLP:conf/icse/SaavedraSM24; DBLP:journals/tse/JiangLLZCNZHBZ23; DBLP:conf/msr/ApplisP23; DBLP:conf/icse/GuanXLLB23; DBLP:conf/issta/XiongX0SWWP0023; DBLP:conf/msr/CsuvikV22; DBLP:journals/softx/PachoulyAK22; DBLP:conf/icse/LiangLSSFD22; DBLP:conf/msr/KeshavarzN22; DBLP:conf/wcre/ZhangYYCYZ22; DBLP:journals/tse/ChenXLGLC22; DBLP:conf/issta/Song0NW0DM22; DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/issre/LiuZDHDM022; DBLP:conf/msr/WendlandSMMHMRF21; DBLP:conf/apsec/AkimovaBDKKMM21; DBLP:journals/stvr/GyimesiVSMBFM21; DBLP:conf/msr/KimKL21; DBLP:conf/splc/NgoNNV21; DBLP:conf/msr/KamienskiPBH21; DBLP:conf/issre/Du0MZ21; DBLP:conf/icse/Makhshari021; DBLP:journals/jss/FerencGGTG20; DBLP:conf/icsm/ZhangXL20; DBLP:conf/msr/KarampatsisS20; DBLP:conf/icst/BuresHA20; DBLP:conf/sigsoft/WidyasariSLQPTT20; DBLP:conf/msr/WangBJS20; DBLP:conf/icse/GarciaF0AXC20; DBLP:conf/promise/VieiraSRG19; DBLP:conf/icse/BentonGZ19; DBLP:conf/icse/DmeiriTWBLDVR19; DBLP:conf/msr/RiganelliMMM19; DBLP:conf/wcre/DelfimUMM19; DBLP:conf/dsa/XuYWA19; DBLP:conf/msr/SahaLLYP18; DBLP:conf/promise/FerencTLSG18; DBLP:conf/issta/ZhangCCXZ18; DBLP:conf/msr/MadeyskiK17; DBLP:conf/icse/TanYYMR17; DBLP:conf/msr/OhiraKYYMLFHIM15; DBLP:journals/tse/GouesHSBDFW15; DBLP:journals/tse/GouesHSBDFW15; DBLP:conf/issta/JustJE14; DBLP:conf/msr/LamkanfiPD13; DBLP:journals/ese/DAmbrosLR12; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/icsm/HoMI0SKNR23; DBLP:journals/tosem/SantanaNAA24; DBLP:journals/infsof/WangBWYGS23; DBLP:journals/jss/WangCHZBZ23; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/kbse/WangWCCY22; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/sigsoft/WangLX0S21; DBLP:conf/kbse/EghbaliP20; DBLP:conf/compsac/GuWL0019; DBLP:conf/kbse/FrancoGR17; DBLP:conf/oopsla/LinKCS17; DBLP:conf/kbse/YeCG23; DBLP:conf/icse/TanDGR18; lu2005bugbench; BegBunch; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:conf/kbse/RomanoLK021; DBLP:conf/msr/AmannNNNM16; DBLP:conf/msr/OliverDAH24; DBLP:conf/sigsoft/GaoDQGW0HZW18; DBLP:conf/sigsoft/Zhang0C0Z19; DBLP:conf/wcre/ReyesGSBM24; DBLP:conf/msr/BeyerGKRR24; DBLP:journals/access/AntalVKMHF24; DBLP:journals/ese/TambonNAKA24; DBLP:journals/pacmpl/DrososSAM024; DBLP:journals/tosem/TufanoWBPWP19; DBLP:conf/kbse/XuZL23) 100
Security (DBLP:conf/kbse/RuanLZL24; DBLP:conf/msr/NiSYZW24; DBLP:conf/pst/TebibAAG24; DBLP:journals/tosem/JiangJWMLZ24; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/ease/WaseemDA0M24; DBLP:conf/kbse/DasAM24; DBLP:conf/qrs/YuW24; DBLP:conf/msr/PramodSTSW24; DBLP:journals/jss/CaiLZZS24; DBLP:conf/icse/WangHGWC024; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/kbse/AnKCYY23; DBLP:conf/raid/0001DACW23; DBLP:conf/secrypt/SenanayakeKAP023; DBLP:journals/compsec/BrustSG23; DBLP:conf/aina/PonteRM23a; DBLP:conf/icse/BhuiyanPVPS23; DBLP:conf/uss/ZhangP0W22; DBLP:conf/msr/BuiSF22; DBLP:conf/prdc/PereiraAV22; DBLP:journals/tse/ChenXLGLC22; DBLP:journals/stvr/GyimesiVSMBFM21; DBLP:conf/icse/PaulTB21a; DBLP:conf/icse/ZhengPLBEYLMS21; DBLP:conf/sigsoft/NikitopoulosDLM21; DBLP:journals/tse/AfroseXRMY23; DBLP:conf/icse/Makhshari021; DBLP:journals/jss/WuZCWM20; DBLP:conf/icsm/ZhangXL20; DBLP:conf/hotsos/MurphyBSR20; DBLP:conf/msr/FanL0N20; DBLP:conf/icse/GarciaF0AXC20; DBLP:conf/msr/RaduN19; DBLP:conf/msr/PontaPSBD19; DBLP:conf/icics/LinXZX19; DBLP:conf/msr/GkortzisMS18; DBLP:conf/promise/MitraR17; DBLP:conf/msr/OhiraKYYMLFHIM15; DBLP:conf/kbse/KuHCL07; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/uss/XuLDDLWPM23; DBLP:journals/jss/WangCHZBZ23; DBLP:conf/fie/SandersWA24; DBLP:conf/icse-apr/YadavW24; DBLP:conf/icse/HuZLYH024; DBLP:conf/issta/WuJPLD0BS23; DBLP:conf/sp/Dolan-GavittHKL16) 48
Performance (DBLP:journals/tosem/LiuMC24; DBLP:conf/icse/CuiD0WSZWYXH0024; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/ease/WaseemDA0M24; DBLP:conf/qrs/YuW24; DBLP:conf/msr/PramodSTSW24; DBLP:conf/icst/KimH23; DBLP:conf/icse/GuanXLLB23; DBLP:conf/msr/KimKL22; DBLP:journals/stvr/GyimesiVSMBFM21; DBLP:conf/issre/Du0MZ21; DBLP:conf/icse/Makhshari021; DBLP:journals/access/SanchezDMS20; DBLP:conf/icsm/ZhangXL20; DBLP:conf/msr/WangBJS20; DBLP:conf/msr/RaduN19; DBLP:conf/issta/ZhangCCXZ18; DBLP:conf/msr/OhiraKYYMLFHIM15; DBLP:conf/msr/ZamanAH12; DBLP:conf/se/SelakovicP17; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/icsm/HoMI0SKNR23; DBLP:conf/msr/AzadIHR23; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/sigsoft/WangLX0S21; DBLP:conf/kbse/FrancoGR17; lu2005bugbench; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:journals/ese/TambonNAKA24; DBLP:journals/pacmpl/DrososSAM024; DBLP:journals/tse/ZhaoXBCL23) 33
Concurrency (DBLP:journals/ese/TimperleyHSDW24; DBLP:conf/asiaccs/LiangYDMHZ23; DBLP:conf/cgo/YuanLLLLX21; DBLP:journals/pacmpl/0001FSK20; DBLP:conf/msr/GaoYJLYZ18; DBLP:conf/kbse/LinMZCZ15; DBLP:conf/hotpar/JalbertPPS11; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/kbse/WangDGGQYW17; lu2005bugbench) 11
Others (DBLP:journals/ese/LiuWWXCWYZ19; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/iwpc/RenL020) 3
The ManyBugs and IntroClass Benchmarks (DBLP:journals/tse/GouesHSBDFW15) are counted as two distinct datasets. This convention applies to all other tables.
This section categorizes datasets based on the types of defects they contain. Across the 151 software defect datasets we survey, we identify four commonly considered defect types: functional bugs, security vulnerabilities, performance bottlenecks, and concurrency issues. In addition, we include a fifth category, others, to account for datasets that do not clearly fall into any of the four primary categories.
Figure 4. Breakdown of types of defects in each application domain.
Table 4 lists the datasets for each category, where a single dataset may appear under one or more categories. Functional defects are the most common type of defects in software defect datasets, with 100 (100/151) datasets in this category. Security vulnerabilities are the second most common type, which are present in 48 (48/151) datasets. Performance defects and concurrency issues are less common, with 33 (33/151) and 11 (11/151) datasets respectively. Figure 4 shows the breakdown of types of bugs for each of the application domains discussed in Section 3.1. We see that functional defects dominate most application domains and other types are also prevalent in specific domains based on the nature of the applications. The rest of this section further discusses these defect types and the datasets that target them. For each defect type, we discuss the characteristics of the defects and highlight representative datasets.
3.2.1. Functional Defects
Functional defects, or functional bugs, refer to errors in software that cause the software to diverge from the intended behavior and to violate its (usually informally provided) specifications. It is the most significant type of software defect, not only in the context of software defect datasets, but also in general software development (Sullivan and Chillarege, 1991; Li et al., 2006). Functional defects impact users in various ways, ranging from minor inconveniences, over incorrect outputs, to application crashes. Figure 4 shows that functional defect datasets dominate most application domains, with the notable exceptions of blockchain and scientific computing.
Functional bugs are the most common focus in software defect datasets, as they form the backbone of software reliability studies. These datasets either provide a general collection of functional bugs across various categories (e.g., Defects4J (DBLP:conf/issta/JustJE14), Bears (DBLP:conf/wcre/DelfimUMM19), Bugs.jar (DBLP:conf/msr/SahaLLYP18), BugSwarm (DBLP:conf/icse/DmeiriTWBLDVR19)) or target a specific subset of functional bugs to facilitate detailed analysis and testing. Examples of such specialized datasets include API misuses (DBLP:conf/kbse/WangDGGQYW17), regular expression defects (Wang et al., 2022), single-statement bugs, (DBLP:conf/msr/KarampatsisS20), type errors (DBLP:conf/wcre/ZhangYYCYZ22), string related defects (DBLP:conf/kbse/EghbaliP20), and numerical bugs (DBLP:conf/kbse/WangWCCY22; DBLP:conf/kbse/FrancoGR17). Overall, functional bug datasets cover a wide variety of topics, making them versatile resources for software defect research.
3.2.2. Security Defects
Security defects, also known as vulnerabilities, are errors in software that make systems vulnerable to attacks or unauthorized access. These defects can compromise the integrity, confidentiality, and availability of a system, leading to severe consequences, such as data breaches, system crashes, or unauthorized access to sensitive information (Landwehr et al., 1994). Security defects are particularly critical in software systems where data protection and privacy are paramount, such as in financial, healthcare, or governmental applications. While these defects can overlap with other types of bugs, their specific focus is on potential exploits that can be leveraged by malicious actors. As security is a major concern in software development, many datasets focus on identifying and analyzing security vulnerabilities to strengthen defenses. Figure 4 shows that datasets of security defects cover 7 out of 11 application domains, ranging from mobile applications to the internet of things. The interest in security defects even surpasses that in functional defects for blockchain-related software, which we attribute to the direct financial impact that vulnerabilities in this domain may have.
Different vulnerability datasets focus on different kinds of defects. One main type of security software defect dataset, e.g., ReposVul (DBLP:conf/icse/WangHGWC024), Big-Vul(DBLP:conf/msr/FanL0N20), and CVEjoin (DBLP:conf/aina/PonteRM23a), focuses on well-defined vulnerabilities, such as those cataloged in standardized databases like CVE (Common Vulnerabilities and Exposures) and categorized by CWE (Common Weakness Enumeration). Another kind of dataset focuses on buffer overflows (DBLP:conf/kbse/KuHCL07), which are critical flaws where programs write memory locations beyond buffer boundaries, and compiler-introduced security bugs (DBLP:conf/uss/XuLDDLWPM23), which arise due to incorrect compiler optimizations. Lastly, cross-language vulnerabilities, which involve multiple programming languages, are captured in datasets like CrossVul (DBLP:conf/sigsoft/NikitopoulosDLM21), providing a basis for multi-language defense techniques.
3.2.3. Performance Defects
Performance defects are software issues that cause performance degradation, leading to inefficient use of resources, increased latency, and longer execution times. These bugs are critical in high-performance computing, web applications, and real-time systems where responsiveness and resource management are essential (Jin et al., 2012). Performance bugs can arise from improper algorithms, inefficient memory usage, excessive computation, or mismanagement of concurrency. These defects can significantly impact user experience and operational costs, making their identification and resolution a priority in software engineering. As shown in Figure 4, datasets of performance issues are prevalent in 9 of 11 application domains, with the exceptions of mobile applications and computational notebooks.
Software defect datasets for performance issues have been constructed for both general applications and specific domains. For example, TANDEM (DBLP:journals/access/SanchezDMS20) offers a taxonomy and dataset of real-world performance bugs, while ECench (DBLP:conf/msr/KimKL22) introduces a benchmark of energy bugs specifically in Ethereum client software, highlighting issues where performance inefficiencies lead to excessive energy consumption. There are also datasets constructed from empirical research of performance issues, such as studies on performance degradation in browser environments (DBLP:conf/msr/ZamanAH12), in JavaScript projects (DBLP:conf/se/SelakovicP17), in database systems (DBLP:conf/icse/CuiD0WSZWYXH0024) and applications (DBLP:journals/tosem/LiuMC24), in high-performance computing (HPC) systems (DBLP:conf/msr/AzadIHR23), in numerical libraries (DBLP:conf/kbse/FrancoGR17), and in cloud systems (DBLP:conf/cloud/GunawiHLPDAELLM14). Such datasets could facilitate future research on pinpointing or mitigating performance issues.
Table 5. Programming languages targeted in defect datasets. One dataset may cover multiple languages.
Language Datasets Count
Java (DBLP:conf/edcc/AndradeLV24; DBLP:conf/msr/LiuHLZCSHM24; DBLP:conf/pst/TebibAAG24; DBLP:journals/tosem/LiuMC24; DBLP:conf/issta/XuG024; DBLP:conf/kbse/DasAM24; DBLP:conf/acl/TianYQCLPWHL0024; DBLP:conf/msr/SilvaSM24; DBLP:conf/icse/WangHGWC024; DBLP:conf/kbse/SongWCCLLWP23; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/icst/LeeKYY24; DBLP:conf/sigsoft/WuLZ024; DBLP:conf/kbse/AvulaVM23; DBLP:conf/secrypt/SenanayakeKAP023; DBLP:journals/tse/JiangLLZCNZHBZ23; DBLP:conf/issta/XiongX0SWWP0023; DBLP:journals/softx/PachoulyAK22; DBLP:conf/msr/KeshavarzN22; DBLP:conf/msr/KimKL22; DBLP:conf/msr/BuiSF22; DBLP:conf/issta/Song0NW0DM22; DBLP:conf/msr/WendlandSMMHMRF21; DBLP:conf/splc/NgoNNV21; DBLP:journals/tse/AfroseXRMY23; DBLP:conf/icse/Makhshari021; DBLP:journals/jss/FerencGGTG20; DBLP:conf/msr/KarampatsisS20; DBLP:conf/icst/BuresHA20; DBLP:journals/pacmpl/0001FSK20; DBLP:conf/msr/WangBJS20; DBLP:conf/icse/DmeiriTWBLDVR19; DBLP:conf/msr/RaduN19; DBLP:conf/msr/RiganelliMMM19; DBLP:conf/wcre/DelfimUMM19; DBLP:conf/msr/SahaLLYP18; DBLP:conf/promise/FerencTLSG18; DBLP:conf/msr/GaoYJLYZ18; DBLP:conf/msr/MadeyskiK17; DBLP:conf/promise/MitraR17; DBLP:journals/ese/LiuWWXCWYZ19; DBLP:conf/kbse/LinMZCZ15; DBLP:conf/issta/JustJE14; DBLP:journals/ese/DAmbrosLR12; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/oopsla/LinKCS17; DBLP:conf/icse/TanDGR18; DBLP:conf/fie/SandersWA24; DBLP:conf/issta/WuJPLD0BS23; DBLP:conf/msr/AmannNNNM16; DBLP:conf/sigsoft/GaoDQGW0HZW18; DBLP:conf/sigsoft/Zhang0C0Z19; DBLP:conf/wcre/ReyesGSBM24; DBLP:journals/tosem/TufanoWBPWP19; DBLP:journals/tse/ZhaoXBCL23; DBLP:conf/kbse/XuZL23) 57
C/C++ (DBLP:journals/ese/TimperleyHSDW24; DBLP:conf/edcc/AndradeLV24; DBLP:conf/msr/LiuHLZCSHM24; DBLP:conf/pst/TebibAAG24; DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:journals/tosem/JiangJWMLZ24; DBLP:conf/icse/CuiD0WSZWYXH0024; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/acl/TianYQCLPWHL0024; DBLP:conf/icse/WangHGWC024; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/kbse/AnKCYY23; DBLP:conf/kbse/AvulaVM23; DBLP:conf/icst/KimH23; DBLP:conf/asiaccs/LiangYDMHZ23; DBLP:conf/raid/0001DACW23; DBLP:conf/icse/GuanXLLB23; DBLP:conf/uss/ZhangP0W22; DBLP:conf/prdc/PereiraAV22; DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/issre/LiuZDHDM022; DBLP:conf/msr/KimKL21; DBLP:conf/icse/PaulTB21a; DBLP:conf/icse/ZhengPLBEYLMS21; DBLP:conf/issre/Du0MZ21; DBLP:conf/icse/Makhshari021; DBLP:conf/msr/FanL0N20; DBLP:conf/icse/GarciaF0AXC20; DBLP:conf/icics/LinXZX19; DBLP:conf/icse/TanYYMR17; DBLP:journals/tse/GouesHSBDFW15; DBLP:journals/tse/GouesHSBDFW15; DBLP:conf/msr/ZamanAH12; DBLP:conf/hotpar/JalbertPPS11; DBLP:conf/kbse/KuHCL07; DBLP:conf/icsm/HoMI0SKNR23; DBLP:conf/uss/XuLDDLWPM23; DBLP:journals/infsof/WangBWYGS23; DBLP:conf/msr/AzadIHR23; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/sigsoft/WangLX0S21; DBLP:conf/compsac/GuWL0019; DBLP:conf/kbse/FrancoGR17; lu2005bugbench; BegBunch; DBLP:conf/icse-apr/YadavW24; DBLP:conf/msr/BeyerGKRR24; DBLP:journals/ese/TambonNAKA24; DBLP:journals/tse/ZhaoXBCL23; DBLP:conf/sp/Dolan-GavittHKL16) 52
Python (DBLP:journals/ese/TimperleyHSDW24; DBLP:conf/msr/LiuHLZCSHM24; DBLP:conf/acl/TianYQCLPWHL0024; DBLP:conf/iclr/JimenezYWYPPN24; DBLP:conf/icse/WangHGWC024; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/sigsoft/WuLZ024; DBLP:conf/kbse/AvulaVM23; DBLP:conf/msr/MahbubSR23; DBLP:journals/jss/ZhaoMLZ23; DBLP:journals/ese/MorovatiNKJ23; DBLP:conf/icse/GuanXLLB23; DBLP:conf/icse/LiangLSSFD22; DBLP:conf/wcre/ZhangYYCYZ22; DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/issre/LiuZDHDM022; DBLP:conf/apsec/AkimovaBDKKMM21; DBLP:conf/msr/KimKL21; DBLP:conf/msr/KamienskiPBH21; DBLP:conf/issre/Du0MZ21; DBLP:conf/icse/Makhshari021; DBLP:conf/sigsoft/WidyasariSLQPTT20; DBLP:conf/msr/WangBJS20; DBLP:conf/icse/DmeiriTWBLDVR19; DBLP:conf/msr/RaduN19; DBLP:conf/issta/ZhangCCXZ18; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/icsm/HoMI0SKNR23; DBLP:journals/tosem/SantanaNAA24; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:conf/kbse/WangWCCY22; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/kbse/FrancoGR17; DBLP:conf/oopsla/LinKCS17; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:journals/access/AntalVKMHF24; DBLP:journals/ese/TambonNAKA24; DBLP:journals/pacmpl/DrososSAM024; DBLP:journals/tse/ZhaoXBCL23) 39
JavaScript (DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/icse/BhuiyanPVPS23; DBLP:conf/msr/CsuvikV22; DBLP:journals/stvr/GyimesiVSMBFM21; DBLP:conf/icse/Makhshari021; DBLP:conf/msr/WangBJS20; DBLP:conf/se/SelakovicP17; DBLP:journals/infsof/WangBWYGS23; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/kbse/EghbaliP20; DBLP:conf/kbse/WangDGGQYW17; DBLP:conf/msr/OliverDAH24; DBLP:conf/kbse/RomanoLK021) 13
Go (DBLP:conf/icse/CuiD0WSZWYXH0024; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/issta/XuG024; DBLP:conf/icse/SaavedraSM24; DBLP:conf/msr/KimKL22; DBLP:conf/cgo/YuanLLLLX21; DBLP:conf/icse/HuZLYH024) 7
C# (DBLP:conf/edcc/AndradeLV24; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/msr/KimKL22; DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/fie/SandersWA24) 5
Rust (DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:conf/qrs/YuW24; DBLP:conf/msr/KimKL22; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/kbse/RomanoLK021) 5
Solidity (DBLP:journals/jss/CaiLZZS24; DBLP:journals/tse/ChenXLGLC22; DBLP:conf/icsm/ZhangXL20; DBLP:journals/jss/WangCHZBZ23) 4
Kotlin (DBLP:conf/pst/TebibAAG24; DBLP:conf/issta/XiongX0SWWP0023; DBLP:conf/icse/BentonGZ19) 3
PHP (DBLP:conf/edcc/AndradeLV24; DBLP:conf/msr/PramodSTSW24) 2
Ruby (DBLP:conf/msr/MadeyskiK17; DBLP:journals/pacmpl/DrososSAM024) 2
Others Fortran (DBLP:conf/kbse/FrancoGR17), F# (DBLP:journals/pacmpl/PaltenghiP22), Groovy (DBLP:conf/icse/BentonGZ19), Haskell (DBLP:conf/msr/ApplisP23), Julia (DBLP:conf/hotsos/MurphyBSR20), Q# (DBLP:journals/pacmpl/PaltenghiP22), Scala (DBLP:journals/pacmpl/0001FSK20) 7
3.2.4. Concurrency Defects
Concurrency defects arise in software when multiple threads or processes execute simultaneously and interact in unintended ways, leading to nondeterministic behavior, crashes, or data corruption. These defects are particularly challenging because they often do not manifest consistently, making them difficult to detect, reproduce, and fix. Concurrency bugs include race conditions, deadlocks, and atomicity violations, which occur when proper synchronization mechanisms fail, or incorrect access to shared resources occurs. Given the complexity of multithreaded systems and the rise of concurrent programming models, identifying and studying concurrency bugs has become critical in ensuring the stability and reliability of modern software systems. As shown in Figure 4, concurrency defects exist in application domains where parallelism and real-time execution are critical, e.g., machine learning, compilers and runtime engines, cloud computing, and autonomous systems.
Several datasets have been developed to address different aspects of concurrency bugs. For C/C++, RaceBench (DBLP:conf/asiaccs/LiangYDMHZ23) provides triggerable and observable race condition bugs to test race detectors, and RADBench (DBLP:conf/hotpar/JalbertPPS11) offers a concurrency bug benchmark suite focused on race conditions, deadlocks, and atomicity violations. For Java, Jbench (DBLP:conf/msr/GaoYJLYZ18) presents a suite of data race bugs specifically designed for concurrency testing, and JaConTeBe (DBLP:conf/kbse/LinMZCZ15) focuses on concurrency bugs in real-world projects, offering manually written test cases for verification. Additionally, there are some datasets of concurrency bugs constructed from empirical studies, e.g., on actor-based concurrency systems (DBLP:journals/pacmpl/0001FSK20), cloud computing systems (DBLP:conf/cloud/GunawiHLPDAELLM14), deep learning compilers (DBLP:conf/sigsoft/ShenM0TCC21), and Node.js (DBLP:conf/kbse/WangDGGQYW17), which offer invaluable insights for future studies.
3.2.5. Other Types of Defects
Some datasets do not fall neatly into the four categories of functional, security, performance, or concurrency bugs. For example, such datasets consider critical bugs that prevent other bugs from being resolved (DBLP:conf/iwpc/RenL020), deployment faults in deep learning-based mobile applications (DBLP:conf/icse/ChenYLCLWL21), or resource leak bugs in Android applications (DBLP:journals/ese/LiuWWXCWYZ19). These datasets address specialized defect types that fall outside the four main categories.
Figure 5. Trend of programming languages in software defect datasets.
3.3. Programming Languages of Defective Software
Another way of categorizing datasets is by the programming language in which the defective code is written. 134/151 of our surveyed defect datasets focus on one or multiple programming languages. These languages include Java, Python, C/C++, JavaScript, Go, and others. Table 5 presents the distribution of programming languages across software defect datasets. Java stands out as the most represented language, with a total of 57 (57/151) datasets, followed by C/C++ (52 datasets, 52/151) and Python (39 datasets, 39/151). The dominance of these languages is likely due to their popularity in software development and due to the focus of the research community on these languages. Other languages, such as F#, Groovy, Haskell, and Julia, have fewer datasets available.
The distribution of languages in defect datasets has evolved over time. Figure 5 shows the trend for the top five programming languages: Java, C/C++, Python, JavaScript, and Go. These five languages also represent the top five most used languages on GitHub (Beuke, 2025), suggesting that dataset creation reflects real-world development trends. Java datasets have increased steadily over the past decade, while Python has gained significant popularity since 2017. JavaScript and Go have experienced peaks within the past five years, and C/C++ datasets have appeared consistently over time. It is important to note the surge in software defect datasets since 2023. While the general increase in publication volume contributes to this rise, another factor is the growing interest in domain- or type-specific defects. The surge is particularly pronounced for C/C++ datasets, with 14 papers in 2023 (a 250% increase from 2022), driven by growing interest in performance bugs (DBLP:journals/tse/ZhaoXBCL23; DBLP:conf/msr/AzadIHR23), security vulnerabilities (DBLP:conf/raid/0001DACW23; DBLP:conf/uss/XuLDDLWPM23), and bugs in emerging domains such as machine learning frameworks (DBLP:conf/icse/GuanXLLB23; DBLP:conf/icsm/HoMI0SKNR23), compilers (DBLP:conf/wcre/WangZRLJ23; DBLP:journals/infsof/WangBWYGS23), and quantum computing (DBLP:conf/qsw/ZhaoWLLZ23). The trend continues in 2024 with 13 C/C++ datasets, suggesting sustained research momentum in infrastructure-level defect analysis. Another contributing factor is the rising demand for large-scale datasets to train and evaluate machine learning and large language models, which has spurred automated bug mining and benchmark construction studies in 2023 (DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/kbse/AnKCYY23; DBLP:journals/tse/JiangLLZCNZHBZ23; DBLP:conf/kbse/SongWCCLLWP23; DBLP:conf/aina/PonteRM23a; DBLP:conf/msr/MahbubSR23; DBLP:conf/raid/0001DACW23; DBLP:conf/icst/KimH23; DBLP:conf/kbse/AvulaVM23; DBLP:conf/kbse/YeCG23). This trend will be further discussed in Section 6.
Figure 6. Programming languages of defects in each domain.
Figure 7. Types of defects for each programming language.
Figure 7 shows the distribution of the most popular programming languages across different application domains. Python and C/C++ dominate machine learning-related defect datasets, and Java and Kotlin cover the major datasets for mobile applications. Compilers and runtime engines have a variety of programming languages, including C/C++, Rust, Python, JavaScript, and Go. Most blockchain-related defects are from Solidity smart contracts, except for one dataset that collects defects from blockchain clients (DBLP:conf/msr/KimKL22), which are implemented in Java, Go, C#, and Rust. In other domains, such as autonomous systems, quantum computing, scientific computing, and computational notebooks, C/C++ and Python are commonly used, whereas datasets for cloud computing, databases, and the internet of things involve Java, Go, C/C++, Python, and JavaScript.
Defect types are not equally distributed across programming languages. Figure 7 shows the distribution of software defect datasets across different programming languages and defect types. Java, C/C++, Python, JavaScript, and Go are common across all defect types, while C#, Rust, Solidity, and PHP cover functional, security, and performance defects only. Kotlin and Ruby have the least coverage, with Kotlin covering only functional and security defects, and Ruby covering only functional and performance defects.
RQ1 Summary Over 60% of the datasets target general-purpose software, while the rest cover eleven specific domains including machine learning, compilers & runtime engines, mobile applications, and blockchain. Machine learning is the most popular specific domain with 14 datasets. Software defect datasets cover functional, security, performance, and concurrency defects. Functional bugs dominate (100/151), followed by security vulnerabilities (48/151), performance issues (33/151), and concurrency defects (11/151). Security and performance datasets are more common in domain-specific contexts such as compilers & runtime engines or blockchain. Java, C/C++, and Python appear in over 75% of all datasets, with JavaScript and Go showing increasing presence recently. Python and C/C++ predominate in machine learning and compilers & runtime engines, while Java and Kotlin dominate mobile applications. Language trends partially align with defect types: Java, C/C++, Python, JavaScript, and Go span all defect types, while other languages associate with specific types.
4. Dataset Construction (RQ2)
Creators of new defect datasets must take various decisions that will impact the quality, quantity, representativeness, and usability of the dataset. In particular, the process of constructing a defect dataset involves selecting an appropriate source and a suitable methodology. There are various sources of defects, such as bug reports from GitHub or Apache Jira, continuous integration systems like Travis-CI or GitHub Actions, and other pre-existing datasets. To extract defects from these sources, existing methodologies range from fully automated, rule-based approaches to semi-automated techniques that incorporate human validation or machine learning models, striving to ensure that the defects collected are representative of real-world scenarios. This section discusses the sources of defects (Section 4.1) and the construction methodologies (Section 4.2) employed by the datasets considered in our study.
4.1. Source of Defects
The source of a defect refers to the origin or platform from which the defect data is collected. As listed in Table 6, there are diverse sources for dataset creators to mine defects from, such as issue reports, version control systems, vulnerability databases, developer Q&A forums, and others. Most of these sources are closely tied to the software development process, as illustrated in Figure 8, especially in software implementation, testing, deployment, and maintenance stages. Other sources, such as coding assignments and defect synthesis, offer a more controlled environment for collecting or creating defects. Among these sources, issue reports and version control systems are the most popular, accounting for 66 (66/151) and 50 (50/151) software defect datasets, respectively. Coding assignments, CI/CD, and defect synthesis are less frequently used, with no more than 8 (8/151) datasets utilizing each. The rest of this section provides a detailed discussion of the primary sources of defects in software defect datasets, highlighting their unique characteristics.
Table 6. Source of defects. One dataset may be based on multiple sources.
Source Datasets Count
Issue Reports (DBLP:journals/ese/TimperleyHSDW24; DBLP:conf/edcc/AndradeLV24; DBLP:journals/tosem/LiuMC24; DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:conf/icse/CuiD0WSZWYXH0024; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/issta/XuG024; DBLP:conf/ease/WaseemDA0M24; DBLP:conf/kbse/DasAM24; DBLP:conf/qrs/YuW24; DBLP:conf/iclr/JimenezYWYPPN24; DBLP:journals/jss/ZhaoMLZ23; DBLP:journals/tse/JiangLLZCNZHBZ23; DBLP:conf/msr/ApplisP23; DBLP:conf/icse/BhuiyanPVPS23; DBLP:conf/icse/GuanXLLB23; DBLP:conf/issta/XiongX0SWWP0023; DBLP:journals/softx/PachoulyAK22; DBLP:conf/icse/LiangLSSFD22; DBLP:conf/msr/KeshavarzN22; DBLP:conf/issre/LiuZDHDM022; DBLP:conf/msr/WendlandSMMHMRF21; DBLP:conf/msr/KimKL21; DBLP:conf/icse/PaulTB21a; DBLP:conf/issre/Du0MZ21; DBLP:conf/icse/Makhshari021; DBLP:journals/jss/FerencGGTG20; DBLP:journals/jss/WuZCWM20; DBLP:conf/icse/GarciaF0AXC20; DBLP:conf/promise/VieiraSRG19; DBLP:conf/msr/RiganelliMMM19; DBLP:conf/msr/PontaPSBD19; DBLP:conf/msr/SahaLLYP18; DBLP:conf/issta/ZhangCCXZ18; DBLP:conf/msr/OhiraKYYMLFHIM15; DBLP:conf/kbse/LinMZCZ15; DBLP:journals/tse/GouesHSBDFW15; DBLP:conf/issta/JustJE14; DBLP:conf/msr/LamkanfiPD13; DBLP:journals/ese/DAmbrosLR12; DBLP:conf/msr/ZamanAH12; DBLP:conf/hotpar/JalbertPPS11; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/se/SelakovicP17; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/icsm/HoMI0SKNR23; DBLP:conf/uss/XuLDDLWPM23; DBLP:journals/infsof/WangBWYGS23; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/kbse/WangWCCY22; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/sigsoft/WangLX0S21; DBLP:conf/iwpc/RenL020; DBLP:conf/kbse/FrancoGR17; DBLP:conf/kbse/WangDGGQYW17; DBLP:conf/icse/TanDGR18; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:conf/kbse/RomanoLK021; DBLP:conf/msr/OliverDAH24; DBLP:conf/sigsoft/GaoDQGW0HZW18; DBLP:conf/msr/BeyerGKRR24; DBLP:journals/access/AntalVKMHF24; DBLP:journals/ese/TambonNAKA24; DBLP:journals/pacmpl/DrososSAM024; DBLP:journals/tse/ZhaoXBCL23) 66
Version Control Systems (DBLP:journals/ese/TimperleyHSDW24; DBLP:conf/pst/TebibAAG24; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/msr/PramodSTSW24; DBLP:conf/icse/WangHGWC024; DBLP:conf/kbse/SongWCCLLWP23; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/icst/LeeKYY24; DBLP:conf/kbse/AnKCYY23; DBLP:conf/kbse/AvulaVM23; DBLP:conf/msr/MahbubSR23; DBLP:journals/ese/MorovatiNKJ23; DBLP:conf/raid/0001DACW23; DBLP:journals/tse/JiangLLZCNZHBZ23; DBLP:conf/msr/ApplisP23; DBLP:conf/msr/CsuvikV22; DBLP:conf/msr/KimKL22; DBLP:conf/wcre/ZhangYYCYZ22; DBLP:conf/msr/BuiSF22; DBLP:conf/issta/Song0NW0DM22; DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/apsec/AkimovaBDKKMM21; DBLP:journals/stvr/GyimesiVSMBFM21; DBLP:conf/cgo/YuanLLLLX21; DBLP:conf/icse/ZhengPLBEYLMS21; DBLP:conf/sigsoft/NikitopoulosDLM21; DBLP:conf/msr/KamienskiPBH21; DBLP:conf/msr/KarampatsisS20; DBLP:conf/sigsoft/WidyasariSLQPTT20; DBLP:conf/msr/WangBJS20; DBLP:conf/icse/BentonGZ19; DBLP:conf/msr/RaduN19; DBLP:conf/dsa/XuYWA19; DBLP:journals/ese/LiuWWXCWYZ19; DBLP:conf/issta/JustJE14; DBLP:journals/tosem/SantanaNAA24; DBLP:journals/jss/WangCHZBZ23; DBLP:conf/msr/AzadIHR23; DBLP:conf/kbse/EghbaliP20; DBLP:conf/compsac/GuWL0019; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:conf/issta/WuJPLD0BS23; DBLP:conf/kbse/RomanoLK021; DBLP:conf/msr/AmannNNNM16; DBLP:conf/msr/BeyerGKRR24; DBLP:journals/access/AntalVKMHF24; DBLP:journals/pacmpl/DrososSAM024; DBLP:journals/tosem/TufanoWBPWP19; DBLP:journals/tse/ZhaoXBCL23; DBLP:conf/kbse/XuZL23) 50
Vulnerability Databases (DBLP:conf/kbse/RuanLZL24; DBLP:conf/msr/NiSYZW24; DBLP:journals/tosem/JiangJWMLZ24; DBLP:conf/icse/WangHGWC024; DBLP:conf/aina/PonteRM23a; DBLP:conf/prdc/PereiraAV22; DBLP:conf/sigsoft/NikitopoulosDLM21; DBLP:conf/msr/FanL0N20; DBLP:conf/icics/LinXZX19; DBLP:conf/msr/GkortzisMS18; DBLP:conf/kbse/KuHCL07; DBLP:conf/kbse/YeCG23; BegBunch; DBLP:conf/icse/HuZLYH024; DBLP:conf/issta/WuJPLD0BS23) 15
Other Pre-Existing Datasets (DBLP:conf/icst/KimH23; DBLP:conf/msr/BuiSF22; DBLP:conf/msr/KamienskiPBH21; DBLP:journals/access/SanchezDMS20; DBLP:conf/icsm/ZhangXL20; DBLP:conf/promise/FerencTLSG18; DBLP:conf/kbse/YeCG23; DBLP:conf/msr/AmannNNNM16; DBLP:conf/msr/OliverDAH24; DBLP:conf/sigsoft/GaoDQGW0HZW18) 10
Developer Q&A Forums (DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:journals/jss/ZhaoMLZ23; DBLP:journals/ese/MorovatiNKJ23; DBLP:journals/tse/ChenXLGLC22; DBLP:journals/pacmpl/0001FSK20; DBLP:conf/issta/ZhangCCXZ18; DBLP:conf/icse/ChenYLCLWL21; DBLP:journals/tosem/SantanaNAA24; DBLP:conf/kbse/WangWCCY22; DBLP:conf/icse/HumbatovaJBR0T20) 10
CI/CD (DBLP:conf/msr/SilvaSM24; DBLP:conf/icse/SaavedraSM24; DBLP:conf/hotsos/MurphyBSR20; DBLP:conf/icse/DmeiriTWBLDVR19; DBLP:conf/wcre/DelfimUMM19; DBLP:conf/msr/MadeyskiK17; DBLP:conf/sigsoft/Zhang0C0Z19; DBLP:conf/wcre/ReyesGSBM24) 8
Defect Synthesis (DBLP:conf/acl/TianYQCLPWHL0024; DBLP:journals/jss/CaiLZZS24; DBLP:conf/asiaccs/LiangYDMHZ23; DBLP:conf/uss/ZhangP0W22; DBLP:conf/splc/NgoNNV21; DBLP:conf/icst/BuresHA20; DBLP:conf/sp/Dolan-GavittHKL16; DBLP:conf/icse-apr/YadavW24) 8
Coding Assignments (DBLP:conf/msr/LiuHLZCSHM24; DBLP:conf/acl/TianYQCLPWHL0024; DBLP:conf/sigsoft/WuLZ024; DBLP:conf/icse/TanYYMR17; DBLP:journals/tse/GouesHSBDFW15; DBLP:conf/oopsla/LinKCS17; DBLP:conf/fie/SandersWA24) 7
Figure 8. Methods for collecting defects (top) from stages in the software development processes (bottom).
4.1.1. Issue Reports
A bug issue report is a documented record of a defect or problem encountered in a software system, typically submitted by developers or users through a bug tracking system, such as GitHub issues or Apache Jira. Issue reports are valuable sources of real-world software defects, providing insight into bugs as documented during the software lifecycle.
Several datasets collect defects directly from GitHub issues. These include gDefects4DL (DBLP:conf/icse/LiangLSSFD22), HasBugs (DBLP:conf/msr/ApplisP23), Andror2 (DBLP:conf/msr/WendlandSMMHMRF21), and Bugs4Q (DBLP:journals/jss/ZhaoMLZ23). In addition to GitHub, Apache Jira serves as another popular bug-tracking system, specifically for the large-scale software projects hosted by the Apache Software Foundation. Datasets like Defects4J (DBLP:conf/issta/JustJE14), ApacheJIT (DBLP:conf/msr/KeshavarzN22), Bugs.jar (DBLP:conf/msr/SahaLLYP18), and DBLP:conf/msr/OhiraKYYMLFHIM15 source defects from Apache Jira, where developers log comprehensive bug reports, including detailed reproduction steps, resolutions, and in many cases, links to the relevant commits that introduce and fix the bugs. Bugzilla is another prominent bug-tracking system that has been leveraged as a defect source by several software defect datasets, including the Eclipse and Mozilla Defect Tracking Dataset (DBLP:conf/msr/LamkanfiPD13), BugHub (DBLP:conf/edcc/AndradeLV24), and CISB (DBLP:conf/uss/XuLDDLWPM23).
These datasets leverage the wealth of defect information available in issue reports, providing robust resources for research on bug prediction, fixing strategies, and software maintenance practices. However, relying exclusively on issue reports introduces several limitations. First, many issue reports lack explicit links to the commits or code regions that caused or fixed the defect, making it difficult to establish accurate traceability between natural-language descriptions and source-code changes (Herzig et al., 2013; Zhou et al., 2012). Second, issue reports are often incomplete or poorly described, missing key environmental or configuration information, which complicates defect reproduction and automated analysis (Bettenburg et al., 2008; Zimmermann et al., 2010). Consequently, datasets built from issue reports typically require extensive manual curation to ensure correctness and reproducibility.
4.1.2. Version Control Systems
The commit histories of version control systems offer a detailed view of software evolution, providing a rich source for collecting software defects. By analyzing commit messages and code changes, researchers can identify bug-fixing commits and their corresponding buggy versions. Several datasets leverage Git commit histories as their primary source of defects. For instance, BugsPHP (DBLP:conf/msr/PramodSTSW24) mines commit pairs from GitHub to capture bug fixes in PHP projects, while Defectors (DBLP:conf/msr/MahbubSR23) focuses on Python, extracting bug-inducing commits through a combination of rule-based and keyword-based approaches. FixJS (DBLP:conf/msr/CsuvikV22) sources bug-fixing in JavaScript commits, grouping them into function-level changes, and Minecraft (DBLP:conf/kbse/AvulaVM23) offers a dataset with precise bug location information mined from Git commit histories. Similarly, D2A (DBLP:conf/icse/ZhengPLBEYLMS21) conducts differential analysis on commit histories, using static analysis tools to identify bug-fixing commits.
Besides GitHub commits, other datasets gather commits from GitLab, such as BugsC++ (DBLP:conf/kbse/AnKCYY23), or utilize the GitHub archive for large-scale multilingual dataset construction, like MBF (DBLP:conf/kbse/ZhongLZGWGL23). These varied sources ensure a broad and diverse collection of bug data across different platforms and programming languages. Collecting defects from version control systems allows dataset creators to capture real-world bugs systematically, but this approach faces challenges. For example, identifying which commits actually introduce or fix defects is difficult, as approaches such as SZZ (Sliwerski et al., 2005, 2005; Kim et al., 2006) often misclassify refactorings or multi-commit fixes (Neto et al., 2018). Additionally, commit messages vary widely in quality, making keyword-based mining unreliable (Tian et al., 2012) and often requiring additional context to fully understand the defect.
4.1.3. Vulnerability Databases & Other Datasets
Another prominent source of software defects is pre-existing datasets, particularly those focused on security vulnerabilities, such as the Common Vulnerabilities and Exposures (CVE) dataset and the National Vulnerability Database (NVD). These datasets provide standardized, well-documented vulnerabilities, which are invaluable for creating benchmarks and conducting empirical studies. For example, CrossVul (DBLP:conf/sigsoft/NikitopoulosDLM21) leverages CVE data to build comprehensive datasets of Java and cross-language vulnerabilities, aimed at studying program repair and vulnerability detection. Similarly, CVEjoin (DBLP:conf/aina/PonteRM23a) consolidates information from CVE and other threat intelligence sources to create a unified dataset for security research. VulinOSS (DBLP:conf/msr/GkortzisMS18) mines vulnerabilities from NVD and CVE to study security issues in open-source projects.
Besides the vulnerability databases mentioned above, some datasets rely on other well-established sources of defects. For example, both PreciseBugCollector (DBLP:conf/kbse/YeCG23) and BugOss (DBLP:conf/icst/KimH23) build defect datasets from OSS-Fuzz (Serebryany, 2017), a continuous fuzzing platform by Google. These datasets, built upon existing structured vulnerability repositories, provide a solid foundation for evaluating and improving security practices, as well as for developing more robust vulnerability detection and repair techniques. However, these datasets are limited to previously mined defects, preventing the inclusion of newly emerging ones. Furthermore, their quality inherently depends on the underlying databases, which can vary significantly across different sources.
4.1.4. Developer Q&A Forums
Q&A platforms, e.g., Stack Exchange and Stack Overflow, provide valuable real-world insights into common bugs faced by developers. These platforms contain a wealth of developer-reported issues, which can be mined and categorized to capture bug symptoms, root causes, and their resolutions. For example, DBLP:journals/ese/MorovatiNKJ23 and DBLP:journals/tse/ChenXLGLC22 extract bugs from Q&A platforms to study defects in machine learning and blockchain systems, respectively. These datasets concentrate on practical, frequently encountered defects accompanied by detailed descriptions and resolutions, making them particularly valuable for empirical research. However, the unstructured nature of data on these platforms poses challenges related to data quality (Ponzanelli et al., 2014), readability (Duijn et al., 2015), and human-factors (Wang et al., 2020), while LLM-based approaches show promise in automating the extraction and structuring of defects from such sources (da Silva et al., 2025).
4.1.5. Continuous Integration & Delivery
A continuous integration and continuous delivery (CI/CD) system records histories of the outcome (e.g., failed or passed) of automated builds and tests after each code change, offering a valuable source of real-world failures and their resolutions. CI/CD is closely tied to version control systems, as it usually monitors code repositories to trigger builds and tests on each commit or pull request. By utilizing fail-pass patterns, several software defect datasets have been built to capture defects that arise during regular development cycles, ensuring their relevance for both research and practice. For instance, BugSwarm (DBLP:conf/icse/DmeiriTWBLDVR19) and Bears (DBLP:conf/wcre/DelfimUMM19) mine CI pipelines from Travis-CI, leveraging fail-pass transitions in the test outcomes to identify and record bug fixes. Similarly, with this model, BugSwarm (DBLP:conf/icse/DmeiriTWBLDVR19; Zhu et al., 2023), GitBug-Java (DBLP:conf/msr/SilvaSM24), and GitBug-Actions (DBLP:conf/icse/SaavedraSM24) also collect a wide range of defects across different software systems from GitHub Actions. These datasets emphasize reproducibility by providing detailed environments and scripts for setting up and replaying failures, making them valuable for execution-based benchmarking and debugging. Moreover, by utilizing the fail-pass pattern from CI/CD systems, these datasets capture real-world defects in an automated, reproducible manner, making them particularly valuable for studies that require large-scale, real-world, and reproducible defect datasets. Nevertheless, collecting defects from CI/CD pipelines inherently excludes defects that do not manifest within these automated processes, potentially overlooking issues that are caught before committing code into the repository or that manifest only in a production environment.
4.1.6. Synthesis
Some datasets generate defects through automated code mutation or bug injection. While this methodology cannot guarantee that the resulting defects are representative of real-world bugs, it provides a controlled and repeatable process for creating benchmarks for research. Specifically, it allows creators to control the location and types of defects, which is not feasible with naturally occurring bugs. DBLP:conf/splc/NgoNNV21 and RaceBench (DBLP:conf/asiaccs/LiangYDMHZ23) introduce synthetic defects, such as variability and concurrency bugs, respectively, to evaluate bug localization and testing tools. FixReverter (DBLP:conf/uss/ZhangP0W22) focuses on bug injection by reverting CVE-related fixes to recreate real-world vulnerabilities for fuzz testing. Similarly, DBLP:conf/icst/BuresHA20 and DBLP:journals/jss/CaiLZZS24 inject faults into open-source projects and smart contracts, respectively, to assess security vulnerability detection. Additionally, with LLMs, DebugBench (DBLP:conf/acl/TianYQCLPWHL0024) injects faulty code into correct LeetCode solutions to create a defect dataset for evaluating the debugging ability of LLMs. These datasets provide a systematic way to study the effects of specific types of bugs and the efficacy of tools designed to detect or repair them.
4.1.7. Coding Assignments
Another source of defects is submissions to coding assignments, programming competitions, and online judging platforms. While these defects may not reflect bugs in complex, real-world software, they often come with detailed descriptions and test cases, making them useful for evaluating bug detection and repair techniques. ConDefects (DBLP:conf/sigsoft/WuLZ024) draws from coding platforms like AtCoder to collect bugs observed in a competitive programming environment. IntroClass (DBLP:journals/tse/GouesHSBDFW15) consists of bugs in student coding assignments and small C programs, providing benchmarks for automated program repair. Codeflaws (DBLP:conf/icse/TanYYMR17) collects defects from competitive programming contests, while QuixBugs (DBLP:conf/oopsla/LinKCS17) is based on the Quixey programming competition, providing a multi-language benchmark for evaluating program repair techniques.
4.2. Construction Methodology
The construction methodology of a software defect dataset refers to the specific process by which defects are collected, and in some cases, validated and categorized. While different datasets can be constructed in radically different ways, some strategies appear repeatedly. Many datasets integrate several distinct strategies at different stages of their construction process, such as combining automated keyword-based search with subsequent manual validation of defects. This means that the construction methodologies outlined in this section are not rigid categories, but instead composable strategies that can be mix-and-matched.
In the following, we separate the construction process into three stages: defect collection, defect validation, and defect categorization. Defect collection involves gathering defects from the sources discussed in Section 4.1. Defect validation is the process of verifying that the collected defects genuinely represent software faults and defect categorization refers to assigning additional classification information to each defect within the dataset. Datasets with fully synthetic defects, as discussed in Section 4.1.6, are excluded from this section, since their construction does not involve defect mining methodologies. However, datasets that obtain code from external sources and subsequently mutate it, such as SBP (DBLP:journals/jss/CaiLZZS24), remain within the scope of this discussion.
4.2.1. Defect Collection
Table 7. Collection methodologies of defects. One dataset may use multiple methodologies.
Methodology Datasets Count
Keyword Matching (DBLP:journals/ese/TimperleyHSDW24; DBLP:journals/tosem/LiuMC24; DBLP:journals/tosem/JiangJWMLZ24; DBLP:conf/icse/CuiD0WSZWYXH0024; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/issta/XuG024; DBLP:conf/ease/WaseemDA0M24; DBLP:conf/kbse/DasAM24; DBLP:conf/msr/PramodSTSW24; DBLP:conf/kbse/SongWCCLLWP23; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/kbse/AvulaVM23; DBLP:conf/msr/MahbubSR23; DBLP:journals/ese/MorovatiNKJ23; DBLP:conf/raid/0001DACW23; DBLP:conf/icse/GuanXLLB23; DBLP:conf/issta/XiongX0SWWP0023; DBLP:conf/msr/CsuvikV22; DBLP:conf/msr/KimKL22; DBLP:conf/wcre/ZhangYYCYZ22; DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/msr/WendlandSMMHMRF21; DBLP:journals/stvr/GyimesiVSMBFM21; DBLP:conf/cgo/YuanLLLLX21; DBLP:conf/msr/KimKL21; DBLP:conf/icse/PaulTB21a; DBLP:conf/msr/KamienskiPBH21; DBLP:conf/icse/Makhshari021; DBLP:journals/access/SanchezDMS20; DBLP:conf/icsm/ZhangXL20; DBLP:conf/msr/KarampatsisS20; DBLP:conf/hotsos/MurphyBSR20; DBLP:journals/pacmpl/0001FSK20; DBLP:conf/msr/WangBJS20; DBLP:conf/icse/GarciaF0AXC20; DBLP:conf/icse/BentonGZ19; DBLP:conf/msr/RaduN19; DBLP:conf/msr/RiganelliMMM19; DBLP:conf/dsa/XuYWA19; DBLP:conf/issta/ZhangCCXZ18; DBLP:journals/ese/LiuWWXCWYZ19; DBLP:conf/kbse/LinMZCZ15; DBLP:journals/ese/DAmbrosLR12; DBLP:conf/msr/ZamanAH12; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/se/SelakovicP17; DBLP:conf/uss/XuLDDLWPM23; DBLP:journals/tosem/SantanaNAA24; DBLP:journals/jss/WangCHZBZ23; DBLP:conf/msr/AzadIHR23; DBLP:conf/kbse/WangWCCY22; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/compsac/GuWL0019; DBLP:conf/kbse/FrancoGR17; DBLP:conf/kbse/WangDGGQYW17; DBLP:conf/icse/TanDGR18; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:conf/kbse/RomanoLK021; DBLP:conf/sigsoft/GaoDQGW0HZW18; DBLP:journals/access/AntalVKMHF24; DBLP:journals/ese/TambonNAKA24; DBLP:journals/pacmpl/DrososSAM024; DBLP:journals/tosem/TufanoWBPWP19; DBLP:journals/tse/ZhaoXBCL23) 64
Mapping (DBLP:conf/kbse/RuanLZL24; DBLP:conf/msr/NiSYZW24; DBLP:journals/tosem/JiangJWMLZ24; DBLP:conf/qrs/YuW24; DBLP:conf/icse/WangHGWC024; DBLP:conf/kbse/SongWCCLLWP23; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/icst/LeeKYY24; DBLP:conf/msr/MahbubSR23; DBLP:conf/raid/0001DACW23; DBLP:journals/tse/JiangLLZCNZHBZ23; DBLP:conf/icse/BhuiyanPVPS23; DBLP:conf/icse/GuanXLLB23; DBLP:conf/issta/XiongX0SWWP0023; DBLP:conf/icse/LiangLSSFD22; DBLP:conf/msr/KeshavarzN22; DBLP:conf/msr/BuiSF22; DBLP:conf/prdc/PereiraAV22; DBLP:conf/apsec/AkimovaBDKKMM21; DBLP:conf/msr/KimKL21; DBLP:conf/icse/PaulTB21a; DBLP:journals/jss/FerencGGTG20; DBLP:journals/jss/WuZCWM20; DBLP:conf/promise/VieiraSRG19; DBLP:conf/msr/RiganelliMMM19; DBLP:conf/msr/PontaPSBD19; DBLP:conf/icics/LinXZX19; DBLP:conf/msr/SahaLLYP18; DBLP:conf/msr/GkortzisMS18; DBLP:conf/kbse/LinMZCZ15; DBLP:conf/issta/JustJE14; DBLP:journals/ese/DAmbrosLR12; DBLP:conf/hotpar/JalbertPPS11; DBLP:conf/icsm/HoMI0SKNR23; DBLP:journals/infsof/WangBWYGS23; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/sigsoft/WangLX0S21; DBLP:conf/iwpc/RenL020; DBLP:conf/kbse/FrancoGR17; DBLP:conf/kbse/WangDGGQYW17; DBLP:conf/icse/TanDGR18; DBLP:conf/issta/WuJPLD0BS23) 42
Manual Collection (DBLP:journals/ese/TimperleyHSDW24; DBLP:journals/tosem/LiuMC24; DBLP:journals/jss/ZhaoMLZ23; DBLP:conf/msr/ApplisP23; DBLP:conf/issre/LiuZDHDM022; DBLP:conf/msr/PontaPSBD19; DBLP:conf/icics/LinXZX19; DBLP:conf/msr/OhiraKYYMLFHIM15; DBLP:conf/hotpar/JalbertPPS11; DBLP:conf/kbse/KuHCL07; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:conf/msr/AmannNNNM16; DBLP:conf/msr/OliverDAH24; DBLP:journals/ese/TambonNAKA24; DBLP:journals/tse/ZhaoXBCL23) 16
External Oracle (DBLP:conf/msr/LiuHLZCSHM24; DBLP:conf/acl/TianYQCLPWHL0024; DBLP:conf/msr/SilvaSM24; DBLP:conf/sigsoft/WuLZ024; DBLP:conf/icst/KimH23; DBLP:conf/icse/SaavedraSM24; DBLP:conf/icse/DmeiriTWBLDVR19; DBLP:conf/wcre/DelfimUMM19; DBLP:conf/icse/TanYYMR17; DBLP:journals/tse/GouesHSBDFW15; DBLP:conf/sigsoft/Zhang0C0Z19; DBLP:conf/wcre/ReyesGSBM24; DBLP:conf/kbse/XuZL23) 13
SZZ Algorithm (Sliwerski et al., 2005) (DBLP:conf/kbse/AvulaVM23; DBLP:conf/msr/MahbubSR23; DBLP:conf/msr/KeshavarzN22; DBLP:conf/icse/PaulTB21a; DBLP:journals/jss/FerencGGTG20; DBLP:conf/msr/KarampatsisS20; DBLP:conf/iwpc/RenL020) 7
Static Analysis (DBLP:conf/icse/WangHGWC024; DBLP:conf/secrypt/SenanayakeKAP023; DBLP:conf/prdc/PereiraAV22; DBLP:conf/icse/ZhengPLBEYLMS21; DBLP:journals/jss/FerencGGTG20; DBLP:conf/msr/KarampatsisS20) 6
Machine Learning (DBLP:conf/icse/WangHGWC024; DBLP:conf/icse/ZhengPLBEYLMS21; DBLP:journals/jss/WuZCWM20) 3
Others (DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:conf/iclr/JimenezYWYPPN24; DBLP:journals/jss/CaiLZZS24; DBLP:journals/compsec/BrustSG23; DBLP:conf/aina/PonteRM23a; DBLP:journals/tse/ChenXLGLC22; DBLP:conf/issta/Song0NW0DM22; DBLP:journals/tse/AfroseXRMY23; DBLP:conf/issre/Du0MZ21; DBLP:conf/sigsoft/WidyasariSLQPTT20; DBLP:conf/msr/GaoYJLYZ18; DBLP:conf/msr/MadeyskiK17; DBLP:journals/tse/GouesHSBDFW15; DBLP:conf/msr/LamkanfiPD13; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:conf/kbse/EghbaliP20; DBLP:conf/msr/BeyerGKRR24) 17
Defect collection refers to the process of gathering defects from one or more mining sources. The particular strategies that dataset creators employ depend on their choice of defect sources and on the intended scope of the dataset. For instance, if a dataset is sourced from CI/CD systems, then collecting defects based on historical test failures is a natural choice. Table 7 lists several common defect collection strategies, showing the diversity of approaches used to construct software defect datasets. Keyword matching and mapping are the most common strategies, with 64 and 42 datasets using them, respectively. Other strategies, such as manual collection, external oracle use, and specialized algorithms (Sliwerski et al., 2005), are less common, but still used by a number of datasets.
Keyword-Based Collection
Many datasets attempt to find defects by scanning for specific keywords (e.g., “bug” or “fix”) in commit messages, issue reports, and other relevant text fields. DBLP:conf/kbse/ZhongLZGWGL23 make for a prototypical example: they examine all GitHub push and pull request events in their target repositories, and consider a bug-fixing event to be any commit or pull request that contains one of the terms “fix,” “repair,” “patch,” “bug,” “issue,” “problem,” or “fault” in the commit message or pull request description. Those keywords are all typical for this kind of approach, built off the assumption that a commit that mentions bugs, repairs, and so on is likely attempting to fix a defect. Keyword-based defect collection has limited precision (Tian et al., 2012), but it is still one of the most common collection approaches observed in this survey, presumably due to its simplicity and its abundance (64 datasets, 64/151) in previous literature.
Mapping
Another common approach in defect collection is to map defects from one source to objects from another source, such as mapping commits from a project’s Git history to related bug reports documented in its issue tracker. This strategy is necessary for datasets that (1) source defects from bug reports or vulnerability databases (Sections 4.1.1 and 4.1.3) and (2) wish to include source-code-level information (Section 5.2). Accordingly, mapping is common, with 42 datasets (42/151) utilizing it. The difficulty of this process can vary based on the mapping source and destination. For instance, mapping GitHub issues to GitHub commits is quite easy, since GitHub tracks commits that reference issues and exposes that relationship in its website and API. In contrast, mapping CVE entries to fixing commits can be more challenging, since many CVE datasets do not reference the commits that fix a vulnerability, and likewise many vulnerability-fixing commits do not reference their associated CVEs. DBLP:journals/tosem/JiangJWMLZ24 outline their process for mapping CVEs to bug-fixing commits in detail, using six different strategies to construct these relationships.
Manual Collection
Some dataset creators forego automatic defect collection, but instead manually collect defects. For instance, DBLP:conf/hotpar/JalbertPPS11 create RADBench by searching through the issue trackers of several projects to find concurrency bugs. An advantage of this approach is that it allows researchers to make qualitative decisions about what defects to include in their datasets, such as only including “vital” issues, which can depend on a variety of factors that are not easy to judge automatically. Conversely, a notable disadvantage is the substantial increase in manual effort imposed on dataset creators. Consequently, it is unsurprising that manually collected datasets typically exhibit a smaller median size compared to automatically collected datasets (298 versus 1158 entries, respectively). Also, while qualitative judgments about a defect can be useful, they are also inherently subjective, which may deter potential dataset users who prioritize objectivity. These disadvantages are probably why this strategy is, with only 16 datasets, relatively uncommon among the surveyed datasets.
External Oracle
Some datasets use historical test failures or continuous integration (CI) failures as external oracles to distinguish buggy from non-buggy software. This should not be confused with datasets that reproduce test failures, which are described in Section 4.2.2. For example, BugSwarm (DBLP:conf/icse/DmeiriTWBLDVR19; Zhu et al., 2023) is created by scanning through the commit history of repositories that use GitHub Actions or Travis-CI, looking for cases where one commit’s CI run fails and a subsequent commit’s run passes. Using historical failures to distinguish bugs from non-bugs allows for better confidence in a dataset’s accuracy, given that failing tests are a tell-tale sign of a bug. However, since this strategy relies on historical test or CI results, it inherently limits the bugs that can appear in a dataset to those that have this historical information available. This naturally limits the defect source of a dataset, which can be an issue if the dataset creators are concerned about dataset size. Moreover, in the case of CI failures, care must be taken to ensure that a given failure was actually caused by a failing test (as opposed to, say, a code style checker).
SZZ-Based Collection
The SZZ algorithm, introduced by Sliwerski et al. (2005) and refined multiple times since (Kim et al., 2006; DBLP:conf/issta/WilliamsS08), attempts to identify the software change that introduced a bug given the change that fixes the bug. This addresses a weakness in some other collection strategies (e.g., keyword search), which is that while they can readily find software changes that fix bugs, they have no way to identify the change(s) that caused those bugs. There are seven datasets in our survey using SZZ or a similar algorithm to mine bug-introducing commits. DBLP:conf/msr/MahbubSR23 demonstrate a typical approach: after finding bug-fixing pull requests in their target repositories, they use SZZ to associate one or more bug-introducing commits to each bug fix.
Static Analysis
There are six datasets in the survey that use static analysis to collect defects. To this end, the dataset creators run existing or newly created static analyses on candidate code. As an example of the former, DBLP:conf/icse/ZhengPLBEYLMS21 create D2A by running the static analyzer Infer on pairs of commits, looking for detected issues that are present in one commit but absent in the next. As an example of the latter, ManySStuBs4J (DBLP:conf/msr/KarampatsisS20) uses a custom static analysis approach that compares the ASTs of candidate buggy and fixed programs and searches for a set of specific bug-fix patterns. An advantage of using static analysis is that, depending on the specific tools or techniques used, one can achieve better precision than strategies such as keyword matching (i.e., fewer non-bugs classified as bugs). The main disadvantage is that relying on static analysis to separate bugs from non-bugs will limit a dataset’s scope to the abilities of the static analysis. Many static analyzers only detect specific classes of bugs (e.g., null pointer bugs), and even more generalist analyzers like Cppcheck will necessarily fail to catch some bugs (Habib and Pradel, 2018).
Machine Learning-Based Collection
There are three datasets in the survey that use machine learning techniques as part of their data collection processes. DBLP:conf/icse/WangHGWC024 input candidate patches into LLMs, asking the models to evaluate whether those patches fix software vulnerabilities. Their dataset also illustrates that extra care must be taken to ensure that LLM output is sound, which they achieve by using LLMs in conjunction with static analysis tools and by including only defects where the two agree. Similarly, DBLP:conf/icse/ZhengPLBEYLMS21 and DBLP:journals/jss/WuZCWM20 use machine learning models to distinguish between defects and non-defects based on semantic information, such as bug reports or commit messages.
4.2.2. Validation
Table 8. Validation methodologies of defects. One dataset may use multiple methodologies.
Methodology Datasets Count
Manual Inspection Full (DBLP:journals/ese/TimperleyHSDW24; DBLP:journals/tosem/LiuMC24; DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:conf/icse/CuiD0WSZWYXH0024; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/issta/XuG024; DBLP:conf/ease/WaseemDA0M24; DBLP:conf/kbse/DasAM24; DBLP:conf/acl/TianYQCLPWHL0024; DBLP:conf/msr/SilvaSM24; DBLP:conf/icst/LeeKYY24; DBLP:journals/jss/ZhaoMLZ23; DBLP:conf/icst/KimH23; DBLP:journals/ese/MorovatiNKJ23; DBLP:conf/msr/ApplisP23; DBLP:conf/icse/GuanXLLB23; DBLP:conf/issta/XiongX0SWWP0023; DBLP:conf/icse/LiangLSSFD22; DBLP:conf/msr/KimKL22; DBLP:conf/wcre/ZhangYYCYZ22; DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/issre/LiuZDHDM022; DBLP:conf/msr/WendlandSMMHMRF21; DBLP:journals/stvr/GyimesiVSMBFM21; DBLP:conf/cgo/YuanLLLLX21; DBLP:conf/icse/PaulTB21a; DBLP:conf/issre/Du0MZ21; DBLP:conf/icsm/ZhangXL20; DBLP:conf/hotsos/MurphyBSR20; DBLP:conf/sigsoft/WidyasariSLQPTT20; DBLP:journals/pacmpl/0001FSK20; DBLP:conf/msr/WangBJS20; DBLP:conf/icse/BentonGZ19; DBLP:conf/msr/RaduN19; DBLP:conf/msr/RiganelliMMM19; DBLP:conf/wcre/DelfimUMM19; DBLP:conf/msr/PontaPSBD19; DBLP:conf/dsa/XuYWA19; DBLP:conf/msr/SahaLLYP18; DBLP:conf/issta/ZhangCCXZ18; DBLP:journals/ese/LiuWWXCWYZ19; DBLP:conf/kbse/LinMZCZ15; DBLP:conf/issta/JustJE14; DBLP:conf/kbse/KuHCL07; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/icsm/HoMI0SKNR23; DBLP:conf/uss/XuLDDLWPM23; DBLP:journals/tosem/SantanaNAA24; DBLP:conf/msr/AzadIHR23; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:conf/kbse/WangWCCY22; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/kbse/EghbaliP20; DBLP:conf/kbse/FrancoGR17; DBLP:conf/kbse/WangDGGQYW17; DBLP:conf/icse/TanDGR18; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:conf/sigsoft/GaoDQGW0HZW18; DBLP:conf/msr/BeyerGKRR24; DBLP:journals/ese/TambonNAKA24; DBLP:journals/tse/ZhaoXBCL23) 62
Partial (DBLP:conf/icse/WangHGWC024; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/kbse/AvulaVM23; DBLP:conf/issta/Song0NW0DM22; DBLP:conf/apsec/AkimovaBDKKMM21; DBLP:conf/msr/KimKL21; DBLP:conf/icse/ZhengPLBEYLMS21; DBLP:conf/icse/Makhshari021; DBLP:journals/jss/FerencGGTG20; DBLP:journals/jss/WuZCWM20; DBLP:journals/ese/DAmbrosLR12; DBLP:conf/msr/ZamanAH12; DBLP:journals/infsof/WangBWYGS23; DBLP:journals/jss/WangCHZBZ23; DBLP:conf/compsac/GuWL0019; DBLP:conf/kbse/RomanoLK021; DBLP:conf/msr/OliverDAH24; DBLP:journals/pacmpl/DrososSAM024; DBLP:journals/tosem/TufanoWBPWP19) 19
Defect Reproduction (DBLP:journals/ese/TimperleyHSDW24; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/acl/TianYQCLPWHL0024; DBLP:conf/iclr/JimenezYWYPPN24; DBLP:conf/msr/PramodSTSW24; DBLP:conf/msr/SilvaSM24; DBLP:conf/icst/LeeKYY24; DBLP:journals/jss/ZhaoMLZ23; DBLP:conf/icst/KimH23; DBLP:conf/icse/SaavedraSM24; DBLP:journals/compsec/BrustSG23; DBLP:journals/tse/JiangLLZCNZHBZ23; DBLP:conf/msr/ApplisP23; DBLP:conf/icse/BhuiyanPVPS23; DBLP:conf/issta/XiongX0SWWP0023; DBLP:conf/wcre/ZhangYYCYZ22; DBLP:conf/msr/BuiSF22; DBLP:conf/issta/Song0NW0DM22; DBLP:conf/msr/WendlandSMMHMRF21; DBLP:journals/stvr/GyimesiVSMBFM21; DBLP:conf/cgo/YuanLLLLX21; DBLP:conf/sigsoft/WidyasariSLQPTT20; DBLP:conf/icse/BentonGZ19; DBLP:conf/icse/DmeiriTWBLDVR19; DBLP:conf/msr/RiganelliMMM19; DBLP:conf/wcre/DelfimUMM19; DBLP:conf/msr/SahaLLYP18; DBLP:conf/issta/ZhangCCXZ18; DBLP:conf/msr/GaoYJLYZ18; DBLP:journals/tse/GouesHSBDFW15; DBLP:conf/issta/JustJE14; DBLP:conf/hotpar/JalbertPPS11; DBLP:conf/se/SelakovicP17; DBLP:conf/uss/XuLDDLWPM23; DBLP:conf/issta/WuJPLD0BS23; DBLP:conf/wcre/ReyesGSBM24; DBLP:journals/access/AntalVKMHF24; DBLP:journals/ese/TambonNAKA24; DBLP:conf/kbse/XuZL23) 39
Dataset validation is about ensuring that the defects included in a dataset are actual defects. This is an important step for making a useful dataset, since a high false positive rate (i.e., many non-bugs classified as bugs) can taint further research that uses a dataset. Even so, not all dataset creators take this step, and there are legitimate reasons for doing so. For instance, a dataset might be so large that validation becomes impractical (e.g., (DBLP:conf/edcc/AndradeLV24; DBLP:conf/msr/CsuvikV22; DBLP:conf/msr/MahbubSR23; DBLP:conf/raid/0001DACW23; DBLP:conf/aina/PonteRM23a)), or a dataset’s data collection approach might already have a low false positive rate (e.g., (DBLP:conf/msr/LiuHLZCSHM24; DBLP:journals/jss/CaiLZZS24; DBLP:conf/kbse/SongWCCLLWP23; DBLP:journals/tse/ChenXLGLC22; DBLP:conf/icse/TanYYMR17; DBLP:conf/sp/Dolan-GavittHKL16)). Table 8 shows common validation methodologies and datasets using them.
Manual Inspection
Some dataset creators, such as DBLP:conf/icst/LeeKYY24, DBLP:conf/msr/WangBJS20, and DBLP:conf/msr/SilvaSM24, opt to manually inspect the code and description of every bug in their datasets (i.e., full manual inspection). This allows researchers to construct a dataset with a low false positive rate and to make qualitative judgments about what bugs should be included, without spending the time and resources needed for manual defect collection. Full manual inspection is quite popular among the datasets surveyed, with 62 (62/151) datasets employing it. The largest (DBLP:conf/acl/TianYQCLPWHL0024) and smallest (DBLP:conf/icst/KimH23) datasets that employ this strategy have 4,253 and 18 entries respectively, with a median of 260 entries.
Conversely, some dataset creators choose to only manually inspect a subset of the defects in their dataset (i.e., partial manual inspection). This is less common than full inspection, with only 19 (19/151) datasets employing it. However, it is attractive for particularly large datasets where full inspection is infeasible, such as D2A (DBLP:conf/icse/ZhengPLBEYLMS21), the largest dataset that uses this approach, which has over 1.3 million entries. The smallest such dataset (DBLP:conf/msr/ZamanAH12) still contains 191 entries, with the median size being 5,565 entries.
Typically, partial inspection has no effect on the actual construction of the dataset, but is performed to demonstrate the accuracy of an automated methodology (DBLP:conf/icse/ZhengPLBEYLMS21). However, there is at least one case where partial inspection is used to filter out defects from a dataset. DBLP:conf/kbse/ZhongLZGWGL23 automatically classify each bug in their dataset into one of multiple fault categories. They then manually analyze 50 bugs from each category to ascertain that category’s accuracy, and only include bugs from categories with over 90% accuracy in the final dataset.
Defect Reproduction
Some dataset creators aim to ensure the accuracy of their datasets by reproducing each defect they collect. Reproduction typically involves running the buggy program in a controlled environment and observing the faulty behavior, often through failing test cases or other verifiable symptoms. While the exact meaning of reproducibility can vary depending on the goals and assumptions of different datasets (Zhu and Rubio-González, 2023), it generally requires the ability to build and run the faulty version of the software.
Among our surveyed datasets, 39 (39/151) involve defect reproduction, and only one dataset attempts to reproduce only a subset of its defects, while the rest aim for full reproduction. This process is often facilitated by sources that provide a rich environment and execution context, such as continuous integration (CI) systems. For example, BugSwarm (DBLP:conf/icse/DmeiriTWBLDVR19) leverages CI configurations and historical build logs to automatically reproduce bugs in Docker-based environments, thus making the automation of reproduction feasible.
However, automation has limitations. Many defects are not covered by existing tests, or their reproduction may require complex setup steps beyond test execution (e.g., user interface-related bugs in AndroR2 (DBLP:conf/msr/WendlandSMMHMRF21)). Still, automated reproduction enables the construction of much larger datasets than manual validation strategies. The largest dataset in this category (DBLP:journals/compsec/BrustSG23) contains approximately 168,000 entries. The median size, 225 entries, is comparable to that of datasets which use full manual inspection.
4.2.3. Categorization
Table 9. Categorization methodologies of defects. One dataset may use multiple methodologies.
Methodology Datasets Count
Manual Categorization (DBLP:journals/ese/TimperleyHSDW24; DBLP:journals/tosem/LiuMC24; DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:journals/tosem/JiangJWMLZ24; DBLP:conf/icse/CuiD0WSZWYXH0024; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/issta/XuG024; DBLP:conf/ease/WaseemDA0M24; DBLP:conf/kbse/DasAM24; DBLP:conf/qrs/YuW24; DBLP:conf/kbse/SongWCCLLWP23; DBLP:journals/jss/ZhaoMLZ23; DBLP:conf/icst/KimH23; DBLP:journals/ese/MorovatiNKJ23; DBLP:conf/msr/ApplisP23; DBLP:conf/icse/GuanXLLB23; DBLP:conf/issta/XiongX0SWWP0023; DBLP:conf/icse/LiangLSSFD22; DBLP:conf/msr/KimKL22; DBLP:conf/wcre/ZhangYYCYZ22; DBLP:journals/tse/ChenXLGLC22; DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/issre/LiuZDHDM022; DBLP:conf/cgo/YuanLLLLX21; DBLP:conf/icse/PaulTB21a; DBLP:conf/msr/KamienskiPBH21; DBLP:conf/issre/Du0MZ21; DBLP:conf/icse/Makhshari021; DBLP:journals/access/SanchezDMS20; DBLP:conf/icsm/ZhangXL20; DBLP:conf/hotsos/MurphyBSR20; DBLP:journals/pacmpl/0001FSK20; DBLP:conf/msr/WangBJS20; DBLP:conf/icse/GarciaF0AXC20; DBLP:conf/msr/RaduN19; DBLP:conf/issta/ZhangCCXZ18; DBLP:conf/msr/GaoYJLYZ18; DBLP:journals/ese/LiuWWXCWYZ19; DBLP:conf/msr/OhiraKYYMLFHIM15; DBLP:conf/kbse/LinMZCZ15; DBLP:journals/tse/GouesHSBDFW15; DBLP:journals/tse/GouesHSBDFW15; DBLP:conf/msr/ZamanAH12; DBLP:conf/hotpar/JalbertPPS11; DBLP:conf/kbse/KuHCL07; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/se/SelakovicP17; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/icsm/HoMI0SKNR23; DBLP:journals/tosem/SantanaNAA24; DBLP:conf/msr/AzadIHR23; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/kbse/WangWCCY22; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/sigsoft/WangLX0S21; DBLP:conf/kbse/EghbaliP20; DBLP:conf/kbse/FrancoGR17; DBLP:conf/kbse/WangDGGQYW17; DBLP:conf/icse/TanDGR18; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:conf/kbse/RomanoLK021; DBLP:conf/msr/AmannNNNM16; DBLP:conf/sigsoft/GaoDQGW0HZW18; DBLP:conf/msr/BeyerGKRR24; DBLP:journals/ese/TambonNAKA24; DBLP:journals/pacmpl/DrososSAM024; DBLP:journals/tse/ZhaoXBCL23) 68
Static Analysis (DBLP:conf/secrypt/SenanayakeKAP023; DBLP:conf/icse/LiangLSSFD22; DBLP:conf/icse/ZhengPLBEYLMS21; DBLP:conf/msr/KamienskiPBH21; DBLP:journals/jss/FerencGGTG20; DBLP:conf/icse/TanYYMR17; DBLP:journals/jss/WangCHZBZ23; DBLP:conf/fie/SandersWA24) 8
Machine Learning (DBLP:conf/kbse/SongWCCLLWP23; DBLP:conf/kbse/AvulaVM23; DBLP:conf/dsa/XuYWA19) 3
Others (DBLP:conf/kbse/RuanLZL24; DBLP:conf/raid/0001DACW23; DBLP:conf/msr/KeshavarzN22; DBLP:conf/msr/BuiSF22; DBLP:conf/prdc/PereiraAV22; DBLP:conf/promise/VieiraSRG19; DBLP:conf/icse/DmeiriTWBLDVR19; DBLP:conf/wcre/DelfimUMM19; DBLP:conf/msr/GkortzisMS18; DBLP:conf/sigsoft/Zhang0C0Z19; DBLP:conf/wcre/ReyesGSBM24) 11
To enhance the utility of their datasets, many dataset creators opt to classify the defects in their datasets along different axes, such as the root causes and symptoms of defects. Defect categorization does not involve adding or removing entries from a dataset, but simply adding metadata to defects already in a dataset. Table 9 shows the categorization strategies for the datasets that use one, which include manual categorization, static analysis, and machine learning.
Manual Defect Categorization
By far the most common way for dataset creators to categorize the defects in their datasets is through manual examination, with 68 (68/151) datasets falling into this category. Similar to manual validation, manual categorization allows researchers to operate reasonably efficiently and flexibly. Indeed, it is not uncommon for datasets to employ both manual validation and manual categorization, (e.g., ECench (DBLP:conf/msr/KimKL22) and ExcePy (DBLP:conf/wcre/ZhangYYCYZ22)), with 48 (48/151) datasets in this group.
Defect categorization need not be entirely automatic or entirely manual. For instance, DBLP:conf/msr/OhiraKYYMLFHIM15 label “surprise” and “dormant” bugs automatically, since the definitions they use are easy to detect, while manually labeling “blocking,” “security,” “performance,” and “breakage” bugs.
Categorization via Static Analysis
Eight datasets in the survey use static analysis to classify defects. For instance, DBLP:conf/secrypt/SenanayakeKAP023 run two static analyzers, Qark and MobSF , on the APKs and source files of Android applications and store the analysis results in their dataset. Categorizing defects based on static analysis has some of the same benefits and drawbacks of static analysis-based collection, namely a fairly high precision, but a limited scope of possible categories and the possibility of false negatives (i.e., bugs with a missing classification).
Categorization via Machine Learning
In total, three of the surveyed datasets classify their entries using machine learning approaches. DBLP:conf/kbse/SongWCCLLWP23 train a neural network to classify bugs given a bug report, fixing commit message, bug-fix location, and bug-inducing location. DBLP:conf/kbse/AvulaVM23 use an off-the-shelf LLM to generate a description of what each bug fix in their dataset does. DBLP:conf/dsa/XuYWA19 use latent Dirichlet allocation (Blei et al., 2003) (a kind of Bayesian network) to extract topic keywords from pull requests and add them to each defect’s metadata.
RQ2 Summary The majority of software defect datasets are sourced from software development processes, with issue reports and version control systems being most common. A smaller number rely on pre-existing datasets, vulnerability databases, developer Q&A forums, coding assignments, and CI/CD systems. Some datasets employ defect synthesis techniques to generate synthetic defects for controlled experimentation. Dataset construction generally involves three stages: defect collection, validation, and categorization. Collection is most often performed through automated mining strategies such as keyword matching (64/151) or mapping (42/151) between issue reports and commits. Fewer datasets rely on manual collection, external oracles, the SZZ algorithm, static analysis, or machine learning techniques. Validation and categorization practices vary widely, though manual inspection remains most common in both stages.
5. Availability and Usability (RQ3)
(a) Availability of artifacts.
(b) Hosting platforms. One dataset may be hosted on multiple platforms.
Figure 9. Availability and hosting platforms of software defect datasets.
The availability and usability of a dataset play a crucial role for its impact on other research and practical applications. Most of the dataset papers in this study provide accompanying artifacts that range from simple download links for datasets to more complex, executable frameworks designed to automate bug reproduction or testing. There are a few instances where the actual datasets are not provided at all, or where the artifacts are not easily accessible or usable. In this section, we examine and discuss the availability and usability of the 151 datasets we survey.
5.1. Availability and Hosting Platforms
Table 10. Availability of artifacts of surveyed datasets.
Availability Datasets Count
Provided Available (DBLP:journals/ese/TimperleyHSDW24; DBLP:conf/edcc/AndradeLV24; DBLP:conf/msr/LiuHLZCSHM24; DBLP:conf/kbse/RuanLZL24; DBLP:conf/msr/NiSYZW24; DBLP:journals/tosem/LiuMC24; DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:journals/tosem/JiangJWMLZ24; DBLP:conf/icse/CuiD0WSZWYXH0024; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/issta/XuG024; DBLP:conf/ease/WaseemDA0M24; DBLP:conf/kbse/DasAM24; DBLP:conf/acl/TianYQCLPWHL0024; DBLP:conf/iclr/JimenezYWYPPN24; DBLP:conf/msr/PramodSTSW24; DBLP:journals/jss/CaiLZZS24; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/kbse/AvulaVM23; DBLP:conf/msr/MahbubSR23; DBLP:journals/jss/ZhaoMLZ23; DBLP:conf/icst/KimH23; DBLP:journals/ese/MorovatiNKJ23; DBLP:conf/asiaccs/LiangYDMHZ23; DBLP:conf/icse/SaavedraSM24; DBLP:conf/raid/0001DACW23; DBLP:conf/secrypt/SenanayakeKAP023; DBLP:journals/compsec/BrustSG23; DBLP:conf/aina/PonteRM23a; DBLP:conf/icse/GuanXLLB23; DBLP:conf/issta/XiongX0SWWP0023; DBLP:conf/msr/CsuvikV22; DBLP:journals/softx/PachoulyAK22; DBLP:conf/icse/LiangLSSFD22; DBLP:conf/msr/KeshavarzN22; DBLP:conf/msr/KimKL22; DBLP:conf/uss/ZhangP0W22; DBLP:conf/wcre/ZhangYYCYZ22; DBLP:conf/prdc/PereiraAV22; DBLP:journals/tse/ChenXLGLC22; DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/msr/WendlandSMMHMRF21; DBLP:conf/apsec/AkimovaBDKKMM21; DBLP:journals/stvr/GyimesiVSMBFM21; DBLP:conf/msr/KimKL21; DBLP:conf/icse/PaulTB21a; DBLP:conf/icse/ZhengPLBEYLMS21; DBLP:conf/sigsoft/NikitopoulosDLM21; DBLP:conf/splc/NgoNNV21; DBLP:journals/tse/AfroseXRMY23; DBLP:conf/msr/KamienskiPBH21; DBLP:conf/icse/Makhshari021; DBLP:journals/jss/FerencGGTG20; DBLP:journals/access/SanchezDMS20; DBLP:journals/jss/WuZCWM20; DBLP:conf/icsm/ZhangXL20; DBLP:conf/msr/KarampatsisS20; DBLP:conf/icst/BuresHA20; DBLP:conf/msr/FanL0N20; DBLP:journals/pacmpl/0001FSK20; DBLP:conf/msr/WangBJS20; DBLP:conf/icse/GarciaF0AXC20; DBLP:conf/promise/VieiraSRG19; DBLP:conf/icse/BentonGZ19; DBLP:conf/msr/RiganelliMMM19; DBLP:conf/msr/PontaPSBD19; DBLP:conf/dsa/XuYWA19; DBLP:conf/icics/LinXZX19; DBLP:conf/msr/SahaLLYP18; DBLP:conf/promise/FerencTLSG18; DBLP:conf/msr/GkortzisMS18; DBLP:conf/issta/ZhangCCXZ18; DBLP:conf/msr/GaoYJLYZ18; DBLP:conf/msr/MadeyskiK17; DBLP:conf/icse/TanYYMR17; DBLP:journals/ese/LiuWWXCWYZ19; DBLP:conf/kbse/LinMZCZ15; DBLP:journals/tse/GouesHSBDFW15; DBLP:journals/tse/GouesHSBDFW15; DBLP:conf/msr/LamkanfiPD13; DBLP:journals/ese/DAmbrosLR12; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/se/SelakovicP17; DBLP:conf/icsm/HoMI0SKNR23; DBLP:conf/uss/XuLDDLWPM23; DBLP:journals/tosem/SantanaNAA24; DBLP:journals/infsof/WangBWYGS23; DBLP:journals/jss/WangCHZBZ23; DBLP:conf/msr/AzadIHR23; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/kbse/WangWCCY22; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/sigsoft/WangLX0S21; DBLP:conf/kbse/EghbaliP20; DBLP:conf/compsac/GuWL0019; DBLP:conf/kbse/FrancoGR17; DBLP:conf/oopsla/LinKCS17; DBLP:conf/kbse/YeCG23; DBLP:conf/icse/TanDGR18; DBLP:conf/fie/SandersWA24; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:conf/icse/HuZLYH024; DBLP:conf/issta/WuJPLD0BS23; DBLP:conf/kbse/RomanoLK021; DBLP:conf/msr/OliverDAH24; DBLP:conf/wcre/ReyesGSBM24; DBLP:conf/msr/BeyerGKRR24; DBLP:journals/access/AntalVKMHF24; DBLP:journals/ese/TambonNAKA24; DBLP:journals/pacmpl/DrososSAM024; DBLP:journals/tosem/TufanoWBPWP19; DBLP:journals/tse/ZhaoXBCL23; DBLP:conf/kbse/XuZL23; DBLP:conf/sp/Dolan-GavittHKL16) 115
Evolving (DBLP:conf/msr/SilvaSM24; DBLP:conf/icse/WangHGWC024; DBLP:conf/icst/LeeKYY24; DBLP:conf/kbse/AnKCYY23; DBLP:conf/sigsoft/WuLZ024; DBLP:journals/tse/JiangLLZCNZHBZ23; DBLP:conf/msr/ApplisP23; DBLP:conf/icse/BhuiyanPVPS23; DBLP:conf/msr/BuiSF22; DBLP:conf/issta/Song0NW0DM22; DBLP:conf/cgo/YuanLLLLX21; DBLP:conf/sigsoft/WidyasariSLQPTT20; DBLP:conf/icse/DmeiriTWBLDVR19; DBLP:conf/msr/RaduN19; DBLP:conf/wcre/DelfimUMM19; DBLP:conf/promise/MitraR17; DBLP:conf/issta/JustJE14; DBLP:conf/msr/AmannNNNM16) 18
No Longer Available (DBLP:conf/kbse/SongWCCLLWP23; DBLP:conf/issre/Du0MZ21; DBLP:conf/hotsos/MurphyBSR20; DBLP:conf/msr/OhiraKYYMLFHIM15; DBLP:conf/kbse/KuHCL07; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/kbse/WangDGGQYW17; BegBunch; DBLP:conf/sigsoft/GaoDQGW0HZW18; DBLP:conf/sigsoft/Zhang0C0Z19) 10
Not Provided (DBLP:conf/pst/TebibAAG24; DBLP:conf/qrs/YuW24; DBLP:conf/issre/LiuZDHDM022; DBLP:conf/msr/ZamanAH12; DBLP:conf/hotpar/JalbertPPS11; DBLP:conf/iwpc/RenL020; lu2005bugbench; DBLP:conf/icse-apr/YadavW24) 8
Ideally, all datasets should be publicly available to facilitate future studies. As shown in Figure 9(a) and Table 10, we found that all but eight datasets were made publicly available at the time of paper publication. In total, 143 (143/150) dataset papers have released their datasets alongside the corresponding paper. For the 143 publicly available datasets, we further examined their availability at the time of our survey and categorized them into the following three groups:
(1) Available means the dataset is still available at the time of our survey, as it was when the corresponding paper was published, but has not been updated. This group accounts for 115/151 of the datasets we survey.
(2) Evolving means the dataset is publicly available, and it has been updated or evolved since the time of paper publication. Such evolution could include adding new defects, deprecating stale defects, or updating the framework or tool to facilitate the use of the dataset. This group accounts for 18/151 of the surveyed datasets.
(3) No Longer Available means the dataset was available at some point in time, but is not available anymore as of the time of our survey. Further investigation shows the reasons include (a) institutional webpage used to host the dataset becoming no longer loadable (DBLP:conf/msr/OhiraKYYMLFHIM15; DBLP:conf/kbse/KuHCL07; DBLP:conf/kbse/WangDGGQYW17; BegBunch), (b) datasets being removed from a hosting service (DBLP:conf/kbse/SongWCCLLWP23; DBLP:conf/issre/Du0MZ21), (c) unresolvable shortened links (DBLP:conf/hotsos/MurphyBSR20), and (d) datasets being lost by accident (DBLP:conf/cloud/GunawiHLPDAELLM14). This group accounts for 10/151 of the surveyed datasets.
Hosting platforms play a significant role in ensuring the ease of access to datasets, as our analysis of the datasets that are no longer available shows that most of them are offline due to discontinued or unstable hosting services. Figure 9(b) shows the distribution of hosting platforms for the available datasets we survey. Note that some datasets are hosted on multiple platforms. The most popular hosting platform is GitHub, which hosts 86 (86/143) available datasets. The second most popular hosting approach is to self-host a dedicated website for the dataset. Other hosting platforms, such as Zenodo, cloud storage services (e.g., Google Drive, OneDrive, Box), are also frequently used, while other platforms such as figshare, Docker Hub, GitLab, Bitbucket, and Mendeley are less common.
5.2. Presentation
There are many ways to present the defects in a dataset, ranging from simply providing the link to the corresponding issue reports to creating complex executable frameworks. The way datasets present defects can impact their usability for both research and practical applications. We categorize the presentation of defects into three levels: (1) meta level, (2) code level, and (3) execution level. Table 11 shows the distribution of these presentation levels across the 133 surveyed datasets that are available or evolving.
Table 11. Presentation of datasets.
Presentation Datasets Count
Meta Level Metadata (DBLP:conf/edcc/AndradeLV24; DBLP:journals/tosem/LiuMC24; DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:journals/tosem/JiangJWMLZ24; DBLP:conf/icse/CuiD0WSZWYXH0024; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/issta/XuG024; DBLP:conf/ease/WaseemDA0M24; DBLP:conf/kbse/DasAM24; DBLP:conf/kbse/AvulaVM23; DBLP:conf/msr/MahbubSR23; DBLP:conf/secrypt/SenanayakeKAP023; DBLP:conf/aina/PonteRM23a; DBLP:conf/icse/GuanXLLB23; DBLP:conf/issta/XiongX0SWWP0023; DBLP:journals/softx/PachoulyAK22; DBLP:conf/icse/LiangLSSFD22; DBLP:conf/msr/KeshavarzN22; DBLP:journals/tse/ChenXLGLC22; DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/msr/KimKL21; DBLP:conf/icse/PaulTB21a; DBLP:conf/sigsoft/NikitopoulosDLM21; DBLP:conf/msr/KamienskiPBH21; DBLP:conf/icse/Makhshari021; DBLP:journals/jss/FerencGGTG20; DBLP:journals/access/SanchezDMS20; DBLP:journals/jss/WuZCWM20; DBLP:conf/msr/FanL0N20; DBLP:journals/pacmpl/0001FSK20; DBLP:conf/msr/WangBJS20; DBLP:conf/icse/GarciaF0AXC20; DBLP:conf/promise/VieiraSRG19; DBLP:conf/msr/RaduN19; DBLP:conf/msr/PontaPSBD19; DBLP:conf/dsa/XuYWA19; DBLP:conf/msr/MadeyskiK17; DBLP:journals/ese/LiuWWXCWYZ19; DBLP:conf/kbse/LinMZCZ15; DBLP:conf/msr/LamkanfiPD13; DBLP:journals/ese/DAmbrosLR12; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/icsm/HoMI0SKNR23; DBLP:conf/uss/XuLDDLWPM23; DBLP:journals/tosem/SantanaNAA24; DBLP:journals/infsof/WangBWYGS23; DBLP:journals/jss/WangCHZBZ23; DBLP:conf/msr/AzadIHR23; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/kbse/WangWCCY22; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/sigsoft/WangLX0S21; DBLP:conf/kbse/EghbaliP20; DBLP:conf/compsac/GuWL0019; DBLP:conf/kbse/FrancoGR17; DBLP:conf/fie/SandersWA24; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:conf/icse/HuZLYH024; DBLP:conf/kbse/RomanoLK021; DBLP:conf/msr/OliverDAH24; DBLP:journals/ese/TambonNAKA24; DBLP:journals/pacmpl/DrososSAM024; DBLP:journals/tosem/TufanoWBPWP19; DBLP:journals/tse/ZhaoXBCL23) 65
Code Level Snippet (DBLP:journals/jss/CaiLZZS24; DBLP:conf/icse/WangHGWC024; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/raid/0001DACW23; DBLP:conf/apsec/AkimovaBDKKMM21; DBLP:conf/icse/ZhengPLBEYLMS21; DBLP:conf/icics/LinXZX19; DBLP:conf/promise/FerencTLSG18) 8
Diff (DBLP:conf/msr/KimKL22; DBLP:conf/uss/ZhangP0W22; DBLP:conf/msr/KarampatsisS20; DBLP:conf/kbse/YeCG23) 4
File (DBLP:journals/jss/ZhaoMLZ23; DBLP:conf/msr/CsuvikV22; DBLP:conf/prdc/PereiraAV22; DBLP:conf/splc/NgoNNV21; DBLP:journals/tse/AfroseXRMY23; DBLP:conf/icsm/ZhangXL20; DBLP:conf/icst/BuresHA20; DBLP:conf/msr/GaoYJLYZ18; DBLP:conf/icse/TanYYMR17; DBLP:journals/tse/GouesHSBDFW15; DBLP:conf/se/SelakovicP17) 11
Execution Level Project (DBLP:journals/ese/MorovatiNKJ23; DBLP:conf/issta/ZhangCCXZ18; DBLP:conf/promise/MitraR17; DBLP:journals/tse/GouesHSBDFW15) 4
Binary (DBLP:journals/compsec/BrustSG23; DBLP:conf/msr/WendlandSMMHMRF21; DBLP:conf/icse/TanDGR18) 3
Framework (DBLP:journals/ese/TimperleyHSDW24; DBLP:conf/msr/LiuHLZCSHM24; DBLP:conf/kbse/RuanLZL24; DBLP:conf/msr/NiSYZW24; DBLP:conf/acl/TianYQCLPWHL0024; DBLP:conf/iclr/JimenezYWYPPN24; DBLP:conf/msr/PramodSTSW24; DBLP:conf/msr/SilvaSM24; DBLP:conf/icst/LeeKYY24; DBLP:conf/kbse/AnKCYY23; DBLP:conf/sigsoft/WuLZ024; DBLP:conf/icst/KimH23; DBLP:conf/asiaccs/LiangYDMHZ23; DBLP:conf/icse/SaavedraSM24; DBLP:journals/tse/JiangLLZCNZHBZ23; DBLP:conf/msr/ApplisP23; DBLP:conf/icse/BhuiyanPVPS23; DBLP:conf/wcre/ZhangYYCYZ22; DBLP:conf/msr/BuiSF22; DBLP:conf/issta/Song0NW0DM22; DBLP:journals/stvr/GyimesiVSMBFM21; DBLP:conf/cgo/YuanLLLLX21; DBLP:conf/sigsoft/WidyasariSLQPTT20; DBLP:conf/icse/BentonGZ19; DBLP:conf/icse/DmeiriTWBLDVR19; DBLP:conf/msr/RiganelliMMM19; DBLP:conf/wcre/DelfimUMM19; DBLP:conf/msr/SahaLLYP18; DBLP:conf/msr/GkortzisMS18; DBLP:conf/issta/JustJE14; DBLP:conf/oopsla/LinKCS17; DBLP:conf/issta/WuJPLD0BS23; DBLP:conf/msr/AmannNNNM16; DBLP:conf/wcre/ReyesGSBM24; DBLP:conf/msr/BeyerGKRR24; DBLP:journals/access/AntalVKMHF24; DBLP:conf/kbse/XuZL23; DBLP:conf/sp/Dolan-GavittHKL16) 38
(1) Meta Level. Meta-level datasets provide the details about defects only in textual form. For example, a table listing all defects with links to corresponding issue reports is considered meta-level. Such metadata may also include the date of creation for an issue report, a severity label, or other annotations. Meta-level datasets are helpful as a reference or resource for analysis but do not include direct means for code analysis or defect reproduction. From the datasets that are still available at the time of this survey, 65 (65/133) are meta-level.
(2) Code Level. Code-level datasets provide isolated code snippets or source files from the defective program, as well as code diffs that show the changes made to the program to fix the defects. For example, a dataset that provides both buggy and fixed versions of source files without providing the entire project is considered code-level. Code-level datasets are helpful for source code analysis, but do not include direct means to run or reproduce the defects. It takes extra effort to integrate (e.g., via git apply) the provided code into a complete project to reproduce the defects. From the datasets that are still available at the time of this survey, 23 (23/133) are code-level.
(3) Execution Level. Execution-level datasets offer more hands-on resources by providing executable artifacts to help reproduce the bugs. For example, a dataset that provides a framework to automatically check out the defective program and reproduce the building and testing process is considered execution-level. Other forms of execution-level datasets include those that provide complete projects with build configurations or datasets that provide executable binaries. Execution-level datasets are particularly useful for practical applications, such as the evaluation of fault localization, test generation, or automated program repair techniques. From the datasets that are still available at the time of this survey, 45 (45/133) are execution-level.
RQ3 Summary Nearly all datasets (143/151) were released publicly at publication. Among them, 115/143 remain available, 18/143 have evolved with updates, and 10/143 are no longer available due to discontinued hosting or link decay. Long-term persistence remains a concern. GitHub dominates hosting, with 86/133 of available datasets. Dedicated websites, Zenodo, and cloud storage (e.g., Google Drive, OneDrive) are also common. Datasets that disappeared were typically hosted on institutional or self-managed servers, indicating that unstable hosting contributes to data loss. Dataset usability varies by presentation level. About 65/133 provide only meta-level information (issue links, descriptions), 23/133 include code-level artifacts (buggy files, diffs), and 45/133 offer execution-level resources (complete projects, reproduction frameworks). Execution-level datasets enable direct evaluation but require additional maintenance.
6. Uses of Software Defect Datasets (RQ4)
Understanding how software defect datasets are used in practice is essential not only for assessing their impact on research and practice, but also for revealing current needs. To this end, we conduct a systematic analysis of research work citing the top-five most-cited [^3] As of January 1, 2025, on Google Scholar. software defect dataset papers among the ones considered in this survey, which are Defects4J (DBLP:conf/issta/JustJE14), Bug Prediction Dataset (DBLP:journals/ese/DAmbrosLR12), Big-Vul (DBLP:conf/msr/FanL0N20), the ManyBugs and IntroClass benchmarks (DBLP:journals/tse/GouesHSBDFW15), and SWE-Bench (DBLP:conf/iclr/JimenezYWYPPN24). All five datasets are general-purpose and not limited to any specific application domain. Four datasets primarily focus on functional defects, while Big-Vul specifically targets security vulnerabilities. In terms of programming languages, Defects4J and the Bug Prediction Dataset collect defects from Java projects; Big-Vul and the ManyBugs and IntroClass benchmarks focus on C/C++ defects; and SWE-Bench, the most recent among them, is in Python.
Methodology
We categorize the uses of software defect datasets by conducting a systematic analysis of papers citing these datasets. Initially, we manually review a subset of citing papers to develop a preliminary categorization. Using an open-coding approach, we continuously refine the categories as we review additional papers. This taxonomy construction process involves over 150 randomly selected citing papers, leading to the following categorization scheme:
  • Empirical Research: Research on existing software defects and existing software engineering techniques, which can be further divided into empirical analyses and empirical evaluations.
    • Empirical Analyses: Papers that analyze existing software defects, such as defect characteristics, root causes, and fix patterns.
    • Empirical Evaluations: Papers that perform retrospective evaluations solely on existing software engineering techniques using software defect datasets.
  • Technical Evaluations: Evaluations of existing or new software engineering techniques using software defect datasets, which can be further divided into detection, testing, and repair.
    • Detection: Papers that use software defect datasets to evaluate bug detection and localization techniques.
    • Testing: Papers that use software defect datasets to evaluate testing techniques, such as test generation and test prioritization.
    • Repair: Papers that use software defect datasets to evaluate automated program repair techniques, such as code repair and issue solving.
  • New Benchmark: Papers that introduce new benchmarks.
  • Others: Papers that cite software defect datasets only as related work but without directly using them.
The “empirical research” and “technical evaluations” categories are not mutually exclusive, as a paper may use a dataset for both purposes. For instance, a paper that empirically evaluates existing program repair tools using established software defect datasets is classified into both empirical research and the repair category of technical evaluations, while a paper that introduces a new repair technique and evaluates it falls solely under technical evaluation.
Based on the above scheme, we use the o3-mini model from OpenAI to automatically categorize papers using their titles and abstracts. The model is instructed with a structured prompt [^4] The full prompt is provided in the supplementary materials (Zhu et al., 2025). that enumerates our taxonomy and labeling rules, including that a paper may have multiple labels, that new benchmark is mutually exclusive with other categories, and that ambiguous cases should be marked as others. Two authors jointly validate a random sample of 100 LLM-labeled papers, independent of the open-coding process, since open coding did not record per-paper labels. All 100 sampled papers have satisfactory classifications: 93 exactly match the human labels, and only 7 show minor discrepancies (e.g., non-contradictory extra categories). For instance, paper (Xia and Zhang, 2024) is labeled as both “Empirical Study” and “Repair” by the model, whereas the human label includes only “Repair”, because the abstract highlights substantial empirical evaluation. Likewise, paper (Ding et al., 2025) is labeled as both “Empirical Study” and “Detection” for the same reason, while the human label includes only “Detection”. We have also inspected the model-provided reasoning summaries, which generally (i) propose candidate categories based on cues in the title and abstract, (ii) apply the taxonomy rules to include or exclude labels, and (iii) resolve overlaps by prioritizing the primary contribution over supporting components.
Figure 10(a) illustrates that empirical research and technical evaluations are the primary use categories for the citing papers analyzed across the five most-cited datasets. Technical evaluations, which can be further divided into evaluations of detection, testing, and repair techniques, demonstrate a growing demand for software defect datasets, as shown in Figure 10(b). This section focuses on these two primary categories to provide a comprehensive overview of the uses of the most-cited software defect datasets. As our analysis specifically focuses on direct dataset utilization, we exclude papers that introduce new benchmarks and the “others” category from the detailed discussion.
6.1. Empirical Research
Software defect datasets can serve as valuable resources for empirical research to understand various aspects of software defects. Such empirical research can be broadly categorized into two groups. On one hand, studies may focus directly on analyzing software defects themselves, typically with regard to their manifestations, symptoms, root causes, and fixes. For example, Sobreira et al. (2018) present an in-depth analysis of 395 patches from Defects4J. By analyzing the patch size, location, and the underlying repair patterns, the study provides insights and future direction for automated program repair, such as highlighting the necessity for repair strategies capable of effectively handling method calls. On the other hand, empirical research can also explore interactions between software defects and existing software engineering techniques. For example, Durieux et al. (2019) conduct an empirical evaluation on eleven automated program repair tools using five software defect datasets. Their study shows that the effectiveness of automated program repair tools varies significantly across different datasets and concludes various causes of failures for the automated program repair. Both types of empirical research significantly benefit from pre-collected, well-structured software defect datasets, as these datasets offer diverse, realistic, and accessible instances of software defects, thus enabling rigorous analysis and reliable comparisons across different studies.
(a) Breakdown of uses.
(b) Trend of techniques evaluated.
Figure 10. Uses of software defect datasets.
6.2. Technical Evaluations
Another primary use of the most-cited software defect datasets is to evaluate existing or new software engineering techniques, such as fault detection/localization (e.g., (Sarhan et al., 2022; An and Yoo, 2022; Schlichtig et al., 2023; Zeng et al., 2022; Yan et al., 2022; Wen et al., 2021; Wang et al., 2021; Yan et al., 2021; Sarhan et al., 2021; Perez et al., 2021; Kim et al., 2021; Nilizadeh et al., 2024; Sun et al., 2024)), test generation/prioritization (e.g., (Zhang et al., 2024d; Jabbar et al., 2023; Rotaru and Vescan, 2023; Chaim et al., 2023; Chen et al., 2023; Almulla and Gay, 2022; Biswas et al., 2022; Gazzola et al., 2022)), and automated program repair (e.g., (Ye et al., 2023; Liu et al., 2023; Tian et al., 2023a; Fan et al., 2023; Chakraborty et al., 2022)). Fault detection/localization aims to identify defects in software in an effective, efficient, and automated manner. Test generation/prioritization is about generating or prioritizing test cases to maximize the likelihood of detecting defects, while minimizing the cost of testing. Both detection and testing techniques have long attracted significant interest from the software engineering, programming languages, and other communities, as such techniques are essential for improving software quality and reliability. The goal of automated program repair (Le Goues et al., 2019) is to fix defective software without human intervention, which is a challenging task that has gained increasing attention in recent years. The evaluation of these techniques is crucial for understanding their effectiveness, efficiency, and limitations, and software defect datasets play a critical role in this process. By providing standardized, realistic, and reproducible instances of software defects, software defect datasets enable researchers and practitioners to evaluate their techniques in a controlled environment.
Figure 10(a) illustrates that detection and repair techniques are the most frequently evaluated. Defects4J is used for evaluating all three types of techniques, and it is also the predominant dataset used for evaluating testing techniques. The Bug Prediction Dataset (DBLP:journals/ese/DAmbrosLR12) and Big-Vul (DBLP:conf/msr/FanL0N20) are primarily used for evaluating detection techniques, while both the ManyBugs and IntroClass benchmarks (DBLP:journals/tse/GouesHSBDFW15) and SWE-Bench (DBLP:conf/iclr/JimenezYWYPPN24) are used more extensively for evaluating repair techniques. Figure 10(b) illustrates the unprecedented growth in the number of proposed software engineering techniques, particularly in the area of automated program repair. This surge in recent years underscores the increasing importance and demand for high-quality software defect datasets.
RQ4 Summary Software defect datasets are primarily utilized in empirical studies and technical evaluations. Empirical work investigates defect characteristics, root causes, and fixes, while technical evaluations emphasize testing, detection, and automated repair. The most-cited datasets are frequently reused in software engineering research. Their adoption indicates a shift toward evaluating automated techniques like automated repair, test generation, and defect detection, highlighting the increasing demand for executable and large-scale datasets suitable for data-driven and LLM-based methods.
7. Discussion and Opportunities for Future Work
In this section, we discuss the challenges and opportunities in creating and maintaining software defect datasets, and provide future directions for enhancing the utility of such datasets for the research community. We focus on challenges related to dataset creation (Section 7.1) and maintenance (Section 7.2).
7.1. Creation of Software Defect Datasets
Scale, Diversity, Quality, and Cost
Though various methodologies for constructing defect datasets have been proposed (Section 4), the balance between scale, diversity, quality, and cost remains a challenge. Intuitively, datasets of high quality tend to be smaller and more costly to create, such as the hand-crafted Defects4J (DBLP:conf/issta/JustJE14). In this sense, further efforts are still needed for fully automated defect mining. However, an important bottleneck for automated mining is the lack of a ground truth that indicates the actual defects present in the codebase. For example, keyword-based mining, the most commonly used technique, shows a relatively low precision of 51.9% and recall of 58.8% (Tian et al., 2012). It would be desirable to establish a ground truth without manual labelling or testing. BugSwarm (DBLP:conf/icse/DmeiriTWBLDVR19) and Bears (DBLP:conf/wcre/DelfimUMM19) have shown that by leveraging the fail-pass pattern in CI/CD pipelines, one can automatically mine defects with high precision. Such techniques may be adopted into the mining processes for other sources, without narrowing to CI/CD build histories. Similarly, ReposVul (DBLP:conf/icse/WangHGWC024) have shown that the combination of LLMs and static analysis can achieve up to 95% precision in defect mining. Such approaches can be further explored to improve the quality of defect datasets while keeping the manual effort low.
Scope Selection
Software defect datasets vary in scope, ranging from general-purpose datasets to domain- or type-specific datasets. Datasets with specialized scopes are desirable for more effective benchmarking of targeted techniques, or for fine-tuning LLMs for specific purposes (Fu et al., 2023). However, as highlighted in Section 3, there is a lack of datasets focusing on specific domains, non-functional defects, and some fast-growing programming languages. For example, despite the widespread adoption of multi-threading capabilities in Python, there is currently no available software defect dataset specifically dedicated to concurrency defects in this language. Future research efforts could prioritize creating datasets that cover more specialized application domains, defect types, or programming languages, to better support the development of targeted software engineering techniques and tools.
Precise Defect Isolation and Annotation
Though tremendous efforts have been made in defect mining, the granularity of defect datasets is usually at the repository-, file-, or method-level, without fully precise defect location information that pinpoints defects, e.g., at the line-level. Ideally, datasets should isolate defects, e.g., by removing any unrelated code changes from a bug-fixing commit. For example, to isolate actual defects from newly added features, refactorings, and other code changes, Defects4J (DBLP:conf/issta/JustJE14) has manually reviewed and filtered the code diffs that fix a given bug. Such a manual review and filtering process is costly and time-consuming, though; thus, it is preferable to have automated techniques to precisely locate defects. Such isolation techniques could be based on refactoring detection (Jiang et al., 2021) or probabilistic delta debugging (DBLP:conf/kbse/SongWCCLLWP23). Another related problem is the annotation of defects, e.g., with the defect’s category. Though some datasets provide such annotations, they are typically manually labelled or inferred from other metadata such as exceptions (DBLP:conf/icse/DmeiriTWBLDVR19). Recently, machine learning and natural language processing techniques have shown promising results for automated defect annotation (DBLP:conf/kbse/SongWCCLLWP23; DBLP:conf/dsa/XuYWA19; Ahmed et al., 2024). More investigation is needed on the effectiveness and efficiency of such techniques, and on how they can be integrated into the process of creating defect datasets.
Dataset Organization
Effectively organizing a dataset is crucial for maximizing its availability and usability. Currently, there is no standardized format for organizing software defect datasets, which can lead to difficulties in dataset access and usage. GitHub is the most popular platform for hosting datasets and tool chains, but it does not ensure long-term dataset availability and is susceptible to data loss or deletion. It is desirable to have datasets archived in a more stable and reliable repository, such as Zenodo or figshare, in addition to GitHub (Association for Computing Machinery (ACM), 2025). Furthermore, the organization of datasets should facilitate future evaluation of new techniques. Section 5.2 shows that half of the datasets provide only textual metadata (i.e., meta-level), which is insufficient for running software defects. Our analysis indicates that, after normalizing citation counts by dataset age (i.e., citations per year), execution-level datasets exhibit higher uptake (mean 53.65, median 12.00) than code-level (mean 21.01, median 12.75) and meta-level datasets (mean 18.07, median 8.50). Future efforts should focus on providing more comprehensive and reproducible software artifacts for defects. At a minimum, we recommend the following artifact reproducibility checklist: (1) a metadata index that records each defect and links it to the corresponding buggy and fixed code versions; (2) an environment descriptor, such as a container image or environment specification; (3) clear build and test instructions with expected outcomes; and (4) a runnable reproduction driver that demonstrates the defect and confirms the fix.
7.2. Maintenance of Software Defect Datasets
Reproducibility
Ensuring the reproducibility of defects in software defect datasets is crucial for researchers to validate their findings and advance existing work. However, achieving reproducibility poses a significant challenge in software defect datasets, regardless of the construction methodology or dataset organization (Zhu and Rubio-González, 2023). Recent work shows that the major reason for defects not being reproducible is inaccessible dependencies, which datasets can mitigate by caching dependencies (Zhu and Rubio-González, 2023; DBLP:conf/icse/SaavedraSM24; DBLP:conf/msr/SilvaSM24). In this direction, more techniques are required to improve the long-term reproducibility of datasets while maintaining the balance between dataset size and usability.
Dataset Evolution
Software defect datasets can benefit from further evolution after initial publication, but most are not updated after release, leading to outdated information. Datasets that do not evolve may cause data leakage when evaluating machine learning-based approaches, especially LLMs, as defects may be in the training corpus (Tian et al., 2023b). Future datasets should support continuous updates by adding new defects while maintaining historical data for longitudinal studies. Continuous mining with automated pipelines (DBLP:conf/icse/DmeiriTWBLDVR19; Zhu et al., 2023) and community-driven efforts (McGuire et al., 2023) show promise for dataset evolution. However, frequent updates complicate comparisons with prior work, as different studies may use distinct dataset versions. Clear versioning, documentation, and stable baselines are crucial for fair and reproducible comparisons. Future research should explore best practices for balancing new defects with historical data maintenance and incentivizing dataset creators to evolve published datasets.
8. Threats to Validity
Selection and implementation bias
Despite our comprehensive search across four major bibliographic databases, we may miss work published outside these databases, or in venues not covered by our selection criteria. Also, excluding preprints and non-peer-reviewed works ensures quality but may miss emerging datasets. Automated scripts in data collection may contain unnoticed errors, and human judgment in screening may introduce subjectivity.
Annotation and interpretation bias
Manual annotation of dataset attributes involves subjective interpretation. Although we measure inter-rater agreement across multiple independent annotations, minor biases may still persist. The LLM-based analysis in Section 6 is manually validated, but some errors may remain due to model limitations.
Construct validity
Our analytical dimensions may omit factors such as community engagement, and differing classification granularity within each dimension may affect cross-study comparability. Our usage analysis prioritizes the most-cited datasets and may not generalize; dataset evolution after publication may also affect the stability of findings.
9. Conclusion
This survey provides a comprehensive overview of the landscape of software defect datasets, including their scope, construction, availability, usability, and practical uses. As fundamental resources to the software engineering, programming languages, computer security, computer systems, and machine learning communities, software defect datasets have seen significant growth in both supply and demand in recent years. This article enables users to identify datasets that best align with their research and practical needs, and it gives future dataset creators guidance on best practices for dataset development and maintenance. Based on the discussions of existing datasets, we also propose several future directions to improve the quality and accessibility of software defect datasets. To facilitate exploration and selection of software defect datasets, we make our artifact publicly available at (Zhu et al., 2025) and provide an interactive website where users can browse and filter datasets based on their specific needs: https://defect-datasets.github.io/.